Nummer 22 März 1998

Mitteilungen der Sprecher . . . . . . . . . . . . . . . . . . . . . 2 Hinweise auf Konferenzen . . . . . . . . . . . . . . . . . . . . . 2 Berichte von Konferenzen . . . . . . . . . . . . . . . . . . . . . 12 Themen und Anwendungen der Computeralgebra . . . . . . 14 M. Clausen, F. Kurth, Audiosignalverarbeitung . . . . . . . . . . . 14 Neues über Systeme und Hardware . . . . . . . . . . . . . . . 17 Maple V Release 5.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Berichte über Arbeitsgruppen . . . . . . . . . . . . . . . . . . 18 Computeralgebra an der RWTH Aachen . . . . . . . . . . . . . . . 18 The FRISCO Project . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Publikationen über Computeralgebra . . . . . . . . . . . . . . 19 Besprechungen zu Büchern der Computeralgebra . . . . . . 20 Grey A., Mezzino M., Pinsky M.A., Introduction to Ordinary Diffe-

rential Equations with Mathematica . . . . . . . . . . . . . . . . 20 Klimek, G., Klimek, M., Discovering Curves and Surfaces with Maple 20 Krawietz, A., Maple V für das Ingenieurstudium . . . . . . . . . . . 21 Vasconcelos, W.V., Computational Methods in Commutative Algebra

and Algebraic Geometry . . . . . . . . . . . . . . . . . . . . . . 22 Lehrveranstaltungen über Computeralgebra im SS 98 … 23 Kurze Mitteilungen . . . . . . . . . . . . . . . . . . . . . . . . . 24

Mitteilungen der Sprecher

Liebe Mitglieder der Fachgruppe Computeralgebra,

die Vorbereitungen auf das Konferenz-Ereignis des Jahres, das

International Symposium on Symbolic and Algebraic Computations, Rostock 1998, 13. August - 15. August 1998

laufen auf Hochtouren, siehe dazu die Ankündigung auf Seite 6. Diese Konferenz, die auf Initiative der Fachgruppe nach Deutschland geholt wurde und nun unter ihrer Verantwortung und der Trägerschaft der GI im Sommer stattfindet, bietet Ihnen die Chance, eine der führenden Konferenzen unseres Fachgebietes ohne allzugroße Reisekosten zu besuchen. Wir laden Sie ein, an diesem Symposium und an den am 12. August stattfindenden Tutorien teilzunehmen. Auch der überwiegende Teil der Fachgruppenleitung wird anwesend sein. Am Rande der Konferenz veranstalten wir eine kurze Informationsveranstaltung für unsere Mitglieder.

Direkt im Anschluß an die ISSAC’98 findet der International Congress of Mathematicans (ICM98) vom 18.-27. August in Berlin statt. Auch dort wird das Thema Computational Mathematics vertreten sein. Weiter wird es eine Special Session on Mathematical Software geben. Näheres zur ICM98 finden Sie auf Seite 7 dieses Rundbriefs.

Als weitere wichtige Konferenz ist das von der Fachgruppe initiierte Symposium

Computeralgebra in Lehre, Ausbildung und Weiterbildung

vom 22. bis 25. April 1998 in Schloß Thurnau bei Bayreuth zu nennen. Informationen dazu finden Sie auf Seite 4.

Auf ihrer Sitzung am 28. Februar in Göttingen hat die Fachgruppenleitung die ersten Vorbereitungen für die turnusgemäß im Herbst stattfindenden Neuwahlen zur Fachgruppenleitung getroffen. Bitte senden Sie Ihre Kandidatenvorschläge bis zum Freitag, 11. September 1998 an den Sprecher. Danach können Vorschläge nur noch bis zum Freitag, 02. Oktober 1998, dem Datum der nächsten Sitzng der Fachgruppenleitung in Künzelsau (Adresse Prof. Werner) angenommen werden, wenn die schriftliche Einverständniserklärung des Kandidaten oder der Kandidatin beigefügt ist.

Johannes Grabmeier B. Heinrich Matzat

Hinweise auf Konferenzen

1. Sixth Rhine Workshop on Computer Algebra Sankt Augustin, 31.3. – 3.4.98

Topics: The workshop intends to cover all aspects of Computer Algebra from theory to applications. As for the previous workshops, we expect that 1/3 of the accepted papers will cover several domains of applications. Purposes: This is the sixth edition of a workshop initiated in Strasbourg in 1988. To avoid competition with wellestablished conferences, the workshop is kept as informal as possible. Its two main purposes are to offer an opportunity to young researchers and newcomers to present their work and to be a regional forum for researchers in this field. Despite this latter goal, the workshop is open worldwide to submissions and attendance. Scientific Committee: Leon Brenig (Brussels), Manuel Bronstein (Zürich), Bruno Buchberger (Linz), Jacques Calmet (Karlsruhe, Chair), Hubert Caprasse (Liège), Alain Carrière (Saint Louis), Jean Della Dora (Grenoble), Vladimir Gerdt (Dubna), Gaston Gonnet (Zürich), Pierre Jost (Strasbourg), Wolfgang Küchlin (Tübingen), Daniel Lazard (Paris), Ton Levelt (Nijmigen), Gerhard Michler (Essen), Teo Mora (Genova), Michael Pohst (Berlin), Fritz Schwarz (Sankt Augustin), Jean Thomann (Strasbourg), Felix Ulmer (Rennes), Volker Weispfenning (Passau). Workshop Chairman: Fritz Schwarz (GMD, Sankt Augustin) Submission: Send two copies of either a full paper or a four pages abstract to the program committee chair. However, electronic mail submissions in Postscript format are encouraged. Submissions will not be formally refereed. They can thus be submitted later to journals for instance. Accepted papers/abstracts will appear in locally printed proceedings (of good quality). These proceedings are intended for attendees only. Updated information will be available at URL http://set.gmd.de/cade/rwca98.html

2. GAMM-Jahrestagung 1998 Bremen, 6.4. – 9.4.98

Die Gesellschaft für Angewandte Mathematik und Mechanik (GAMM) lädt zu ihrer wissenschaftlichen Jahrestagung an der Universität Bremen vom 6. bis 9. April 1998 ein. Die ordentliche Mitgliederversammlung der GAMM findet am Mittwoch, dem 8. April 1998 um 11:30 Uhr, im Großen Hörsaal der Universität statt.

Tagungsleitung:

H.J. Rath, D. Hinrichsen, M. Wanschura, B. Schuldt, L. Arnold, A. Bunse-Gerstner, M. Dreyer, Ch. Egbers, R. Kienzler, H.C. Kuhlmann, H.-O. Peitgen, P.H. Richter, alle an der Universität Bremen.

In den Sektionen finden Kurzvorträge über aktuelle Themen aus den Gebieten der angewandten Mathematik und der Mechanik statt. Auch Anwendungen der Mathematik auf ingenieurwissenschaftliche Probleme, die nicht der Mechanik entstammen, sind mögliche Themen. Die Vortragszeit beträgt 15 Minuten, weitere 5 Minuten stehen für die Diskussion zur Verfügung. Vortragssprachen sind Deutsch, Englisch und Französisch. Im Rahmen der Sektionen gibt es ein bis zwei Vorträge mit einer Vortragszeit von 30 Minuten. In diesen Vorträgen soll über wichtige neuere Entwicklungen der betreffenden Gebiete, auch in Form von Übersichten, berichtet werden. Die Sektionssitzungen finden an den Nachmittagen statt und beginnen jeweils um 14:00 Uhr. Die Sektionen entnehmen Sie bitte der folgenden Aufstellung:

1: Lineare und nichtlineare Schwingungen, 2: Stabilität und Regelung, 3: Mehrkörpersysteme und Kinematik, 4: Elastische und viskoeleastische Stoffe und Systeme, 5: Plastizität: Stoffgesetze und Anwendungen, 6: Schädigungsund Bruchmechanismen, 7: Numerische Mechanik, 8: Experimentelle Methoden und Identifikation, 9: Reibungsfreie Strömungen, Gasdynamik, 10: Reibungsbehaftete Strömungen, Turbulenz, 11: Wärme- und Stoffübertragung, Konvektionsströmungen, 12: Mehrphasenströmungen, einschl. Tropfen und Blasen, 13: Wellen, 14: Angewandte Analysis, 15: Mathematische Methoden in den Ingenieur- und Naturwissenschaften, 16: Computeralgebra und -analysis (Sektionsleitung: K. G. Roesner, TH Darmstadt und J. L. Grabmeier, IBM Heidelberg), 17: Angewandte Stochastik, Operations Research 18: Optimierung, 19: Numerische Analysis, 20: Numerische Behandlung von gewöhnlichen Differentialgleichungen und Algebrodifferentialgleichungen, 21: Numerische Behandlung von partiellen Differentialgleichungen 22: Kontroll- und Systemtheorie, 23: Finanzmodellierung, 24: Raumtransportsysteme.

Weitere Informationen finden Sie im Einladungsheft, das im September an die GAMM-Mitglieder mit dem GAMM-

Rundbrief versandt wird und im Internet unter http://www.zarm.uni-bremen.de/gamm98/gamm98d.htm zu finden

ist.

Als Email-Adresse verwenden Sie bitte: gamm98@zarm.uni-bremen.de

3. Conway Fest – JHC60 Cambridge, England, 16.4. – 17.4.98

A conference to mark the 60th birthday of John Conway will be held at the Newton Institute, Cambridge, on 16th and 17th April 1998. The event will be partially supported by the London Mathematical Society and Gonville and Caius College.

Talks: N.J.A. Sloane (AT&T Labs-Research), Recent results on packing spheres in n-dimensional space, D. Zagier (Bonn), The Dedekind eta-function, a mysterious identity, and knot invariants, E. Berlekamp (Berkeley), Sums of Nx2 Amazons positions, S. Kochen (Princeton), Geometry, Logic and Quantum Theory, B. Fischer (Bielefeld), Remarks on sporadic simple groups, R.E. Borcherds (Cambridge), Hyperbolic reflection groups (after Conway and Allcock), L. Kauffman (UIC), Conway’s Knots and Their Ramifications, J.H. Conway (Princeton), The Fifteen theorem for quadratic forms.

4. Workshop on Computer Algebra in Scientific Computing Sankt-Petersburg, Russia, 20.4. – 24.4.98

Topics:

The workshop will cover all aspects of the use of computer algebra and symbolic computation in scientific computing including both theory and applications. Among topics of particular interest they are:

numerical simulation with computer algebra systems, parallel symbolic-numeric computations, symbolic-numeric interface, symplectic integration, construction of approximate solutions of differential equations and dynamical systems, celestial mechanics and general relativity, algebraic methods for nonlinear equations and inequalities, computer algebra methods in pure mathematics, computational group theory, applications to the theory of error-correcting codes, problem-solving environments for partial differential equations.

Workshop co-chairs: Vladimir Gerdt (Dubna), Ernst Mayr (Munich).

Program Committee:

Evgenii Vorozhtsov (Novosibirsk, co-chair), Christoph Zenger (Munich, co-chair), Victor Edneral (Moscow), Victor Ganzha (Munich), Gaston Gonnet (Zurich), Gert-Martin Greuel (Kaiserslautern), Stanly Steinberg (U.S.A.), Yuri Matiyasevich (St. Petersburg), Vitaly Mysovskikh (St. Petersburg), Alexander Skopin (St. Petersburg), Nikolay Vassiliev (St. Petersburg), Volker Weispfenning (Passau).

Local Organizing Committee:

Alexander Skopin (co-chair), Nikolay Vassiliev (co-chair), Sergey Baranov, Vitaly Mysovskikh, Elena Novikova (secretary).

Further Information: http://www.pdmi.ras.ru/EIMI/comalg.html,
comalg@euler.pdmi.ras.ru, casc98@euler.pdmi.ras.ru,
vasiliev@ita.runnet.ru (Nikolay Vassiliev)

5. Computeralgebra in Lehre, Ausbildung und Weiterbildung Bayreuth, 22.4.–25.4.98

Die Fachgruppe Computeralgebra (FG CA) der DMV (Deutsche Mathematiker–Vereinigung), GAMM (Gesellschaft für Angewandte Mathematik und Mechanik) und der GI (Gesellschaft für Informatik) organisiert diese Tagung gemeinsam mit der GDM (Gesellschaft für Didaktik der Mathematik), der MNU (Verein zur Förderung des mathematischen und naturwissenschaftlichen Unterrichts) und der Fachgruppe Didaktik der Mathematik der DMV (FG DM).

Ziel ist es, zu diesem Thema einen breiten Austausch zwischen den Kultusministerien, den für die Fortentwicklung der curricularen Lehrpläne zuständigen Instituten und den Experten aus Wissenschaft und Lehre der organisierenden Gesellschaften zu initiieren. Wir erhoffen uns insbesondere Berichte über die in den einzelnen Bundesländern stattfindenden Lehrversuche und Pläne.

Verantwortlich ist das Organisations- und Programmkomitee in Zusammenarbeit mit der Fachgruppenleitung Computeralgebra. Es besteht aus

• Prof. Dr. A. Kerber, Bayreuth (FG CA, Leitung und lokale Organisation),

• Prof. Dr. Wolfram Koepf, Leipzig (FG CA),

• Dr. Günther Schmidt, Mainz (MNU),

• Prof. Dr. Günter Törner, Duisburg, (FG DM) und

• Prof. Dr. Hans-Georg Weigand, Gießen (GDM).

Wegen des auf ca. 45 Teilnehmer beschränkten Umfangs bitten wir um frühzeitige Anmeldung.

Vorträge:

D. Pohlmann (Ministerium Schleswig- Holstein, Elmshorn), H. Heugl (Wien), Lehren und Lernen mit Computeralgebra-Systemen – Didaktische Konzepte und Erfahrungen aus den österreichischen CAS-Projekten, B.H. Matzat (Interdisz. Zentrum f. wiss. Rechnen, Univ. Heidelberg), F. Postel (Mathem./Informatik Paderborn), Ein Autorensystem zur Erstellung multimedialer mathemischer Unterrichtseinheiten, W. Lichtenberg (Landesinstitut für Lehrerfortbildung von an Gymnasien, Sachsen-Anhalt), Stand von Erprobungen beim Einsatz von Computeralgebra-Systemen im Unterricht Mathematik, P. Rasfeld (Inst. f. Didaktik der Mathematik, Dortmund), H. Kilian (Univ. Dortmund), Einsatz von CAS-Systemen zur Konstruktion und Kontrolle von Beweisen durch Studenten und Schüler, J. Wagner (Comenius-Inst. Radebeul), W. Zappe (Ilmenau), Der TI-92 im Mathematikunterricht – Erfahrungen eines Lehrers aus Thüringen, S. Griebel (Texas Instruments - E&P Freising), Überblick über die Zulassung graphischer und symbolischer Taschenrechner in Deutschland und Europa, H.-W. Henn ( Staatl. Seminar für Schulpädagogik (Gymn) Karlsruhe), Das Pilotprojekt “Mobiles Klassenzimmer”in Baden-Württemberg, W. Herget (Univ. Halle-Wittenberg), Kleine Schritte sind besser als keine – und besser als zu große, C. Richard (Institut f. Theor. Physik Universität Tübingen), Analysis Alive - Ein interaktiver Mathematik-Kurs, U. Klein (Lehrstuhl Didaktik für Mathematik Aachen), V. Reineke (Niedersächsisches Kultusministerium Hannover), F. Schumann (Math-College Hannover), Zuordnungen – Grundbegriffe für einen erweiterten Funktionsbegriff in Sek.I, W. Neidhardt, A. Wassermann (Lst Mathematik und ihre Didaktik Bayreuth), Demonstration – Geonet – Dynamische Geometriesoftware mit JAVA, W. Koepf (FB IMN, HTWK Leipzig), P. Dräxler (Fakultät f. Mathematik Universität Bielefeld), CREP – Darstellungstheorie mit Computeralgebra in Forschung und Lehre.

6. ECCAD’98 – Fifth Annual East Coast Computer Algebra Day Annapolis, Maryland, USA, 25.4.98

The fifth annual East Coast Computer Algebra Day will be held at the U.S. Naval Academy in the beautiful city of Annapolis, Maryland, on April 25, 1998.

The previous such meetings were held at: 1. Drexel University (Philadelphia, Pennsylvania), 2. University of Delaware (Newark, Delaware), 3. IBM Watson Laboratories (Yorktown Heights, New York), 4. Northeastern University (Boston, Massachusetts),

The program consists of: (a) Invited talks (b) Poster sessions (c) Math software vendors (d) Informal gatherings, beverages, and lunch. (and dinner for those who wish to participate.)

Invited talks: Rob Corless, University of Western Ontario, Canada (Title to be announced.),

Scott Mccallum, Macquarie University, Australia, on ”Factors of iterated resultants and discriminants.”

Erich Kaltofen, Rensselaer Polytechnic Institute, on ”Challenges of symbolic computation: my favorite open problems.”

Posters:

To present a poster: Please email to George Nakos at gcn@nadn.navy.mil:

(1) Name, Address, Affiliation, E-mail, (2) Poster title, (3) Poster abstract (not to exceed 250 words. In ASCII, or ASCII LateX.).

Topics for posters include: Symbolic, algebraic, and symbolic-numerical algorithms. Computer algebra systems, software systems, experience with implementations of significant algorithms. Applying symbolic, algebraic, and symbolicnumerical algorithms to problems in the sciences, engineering, economics, and other areas.

In addition to the usual research posters we encourage posters on the use of Computer Algebra Systems in the Classroom.

Abstracts of the posters will be published in the SIGSAM Bulletin, and a version of posters to be displayed on the Web is also solicited.

Registration: Registration is free. We encourage all participants to register by emailing to Peter Turner at prt@nadn.navy.mil, Your: Name, Address, Affiliation, E-mail

Vendors:

Vendor displays and software demonstrations are also anticipated. If you are interested in such a display or demon-

stration, please send e-mail to prt@nadn.navy.mil or gcn@nadn.navy.mil
Further information: http://turnerpr.sma.usna.navy.mil/ECCAD98/

7. Tagung zum Unterrichtseinsatz des symbolischen Taschenrechners TI-92 Münster, 2.6. – 5.6.98

Vom 2.6.1998 bis zum 5.6.1998 findet an der Westfälischen Wilhelms-Universität Münster eine Tagung zum Unterrichtseinsatz des symbolischen Taschenrechners TI-92 statt. In Vorträgen und Workshops werden methodisch-didaktische Konzepte des Unterrichtens mit Computeralgebra-Systemen vorgestellt und praktisch erprobt.

Tagungsleitung: Westfälische Wilhelms-Universität Münster Zentrale Koordination Lehrerausbildung StR’ Bärbel Barzel, Düsseldorf, Dr. Detlef Berntzen, Münster.

Adressatenkreis: Lehrerinnen und Lehrer weiterführender Schulen, Fachleiter/innen, Referendar/innen, Fachdidaktiker/innen Mathematik

Teilnehmerbeitrag: 100 DM (inkl. Empfang und Sozialprogramm)

Detaillierte Infos und Tagungsplan unter: http://www.uni-muenster.de/Lehrerausbildung/ti92tag.htm oder

direkt bei: Westfälische Wilhems-Universität Münster Zentrale Koordination Lehrerausbildung, Dr. Detlef Berntzen, Email: berntz@uni-muenster.de, Telefon: 0 25 1/ 51 03 82 1, Prinzipalmarkt 38, 48143 Münster.

Letzte Meldung: folgende Referenten haben ihr Erscheinen zugesagt: Prof. Dr. Bert Waits, Ohio Stat University, USA, Prof. Dr. John Berry, University of Plymouth, UK, Paul Drijvers, Freudenthal-Institut, Utrecht, NL, Günter Schmidt, Stromberg, D.

8. Worldwide Mathematica Conference Chicago, USA, 18.6. – 21.6.98

This year is Mathematica’s 10th anniversary, and to celebrate we’re hosting the Worldwide Mathematica Conference in Chicago on June 18-21, 1998. Since you’re a registered Mathematica user, we hope that you’ll be able to join us at this exciting and unique event.

The conference will bring Mathematica users from around the world together to discover and explore the latest in Mathematica technology. The program will include distinguished keynote speakers, tutorials at all levels, and presentations and demonstrations of Mathematica in action–in research, in education, and in technical communication. There will also be opportunities for Mathematica users such as yourself to present work in a variety of formats.

To cap the celebration of 10 years of Mathematica, there will be an exclusive dinner and reception at Chicago’s Museum of Science and Industry.

The registration fee for academic users is 595(495 if you register by March 31). Discounts are available to Premier and Regular Service subscribers. The fee for Premier Service subscribers is 495(395 if you register by March 31). The fee for Regular Service subscribers is 520(420 if you register by March 31).

For more conference information and an online registration form, visit our web site at

http://www.wolfram.com/conference98. If you have questions, send email to
conference98@wolfram.com or call us at 217-398-0700.

9. IMACS - ACA’98 The Fourth International IMACS Conference on Applications of Computer Algebra Prag, August 9.8. – 11.8.98

Scope of the Conference: The meeting will focus on actual or possible applications of nontrivial computer algebra techniques to other fields and substantial interactions of computer algebra with other fields.

Meeting Format: The meeting will be run in the standard IMACS format where individuals are invited to organize a special session. Individuals can propose a special session by contacting the program chairs. All paper submissions must be directed to an organizer of an appropriate special session, which will be listed on the web sides.

Preliminary List of Proposed Sessions: Mechanic, ODEs, Problem Solving Environments, High Energy Physics, PDEs and Symmetries, Symbolic-numeric Interfaces, General Relativity, Hamiltonian Systems, Seminumerical Methods, Robotics, Geometry, Automatic Differentiation, Control Theory, Grand Challenges, Interval Arithmetic, Industrial Applications, Software Integration, Education.

Organization: General Chair: Richard Liska (liska@siduri.fjfi.cvut.cz) Program Chairs: Victor Edneral (edneral@theory.npi.msu.su),

Wolfgang Küchlin (kuechlin@informatik.uni-tuebingen.de)

Local arrangements: Jiri Limpouch, Stepan Müller, Milan Sinor, Faculty of Nuclear Sciences and Physical Engineering, Czech Technical University.

More Information: http://www-troja.fjfi.cvut.cz/aca98 or send e-mail to aca98@siduri.fjfi.cvut.cz.

10. ISSAC’98 – International Symposium on Symbolic and Algebraic Computation Rostock, 13.8. – 15.8.98

ISSAC is a yearly international symposium that provides an opportunity to learn of new developments and to present original research results in all areas of symbolic mathematical computation.

Topics of the meeting include, but are not limited to: Algorithmic mathematics Algebraic, symbolic, and symbolic-numeric algorithms including: simplification, polynomial and rational function manipulations, algebraic equations, summation, integration, linear algebra and matrix computations, number theory, ODE/PDE, complex computation, group computations, and geometric computing; Computer science Theoretical and practical problems in symbolic mathematical computation including: computer algebra systems, problem solving environments, programming languages and libraries for symbolic computation, user interfaces, data structures, software architectures, parallel/distributed computing, mapping algorithms to architectures, concrete analysis and benchmarking, complexity of computer algebra algorithms, artifical intelligence techniques, automatic differentiation and code generation, mathematical data exchange protocols; Applications Problem treatments incorporating algebraic, symbolic or symbolic-numeric computation in an essential or novel way, including engineering, economics and finance, physical and biological sciences, computer science, logic, mathematics, statistics, and use in education.

Conference activities The planned activities include invited presentations, research and survey papers, poster sessions, tutorial courses, vendor exhibits, and software demonstrations. Proceedings will be distributed at the symposium. Proposals for workshops, tutorial courses, demonstrations, panel discussions or related activities are welcomed. Usergroups, editorial boards or other associations desiring meeting space during the course of the symposium are encouraged to contact the conference organizers.

Conference Location ISSAC’98 will be held in the historic buildings of the University of Rostock, Germany. Accomodations will be available at nearby hotels. The old “Hansestadt” Rostock is situated close to the Baltic Sea and about 200 km away from Berlin. Participants have the occasion to attend afterwards the International Congress of Mathematicians ICM’98 held August 18-27 in Berlin.

General Chair: Volker Weispfenning, FMI, Universität Passau, D-94030 Passau, Germany, issac98@alice.fmi.uni-passau.de (+49) 851-509-3120 phone, (+49) 851-509-3122 fax.

Program Committee Chair: Barry Trager, IBM T.J. Watson Research Center, P.O.Box 218, Yorktown Heights,

     N.Y. 10598, USA. bmt@watson.ibm.com

Program Committee Vice-Chair: Robert M. Corless, Dept. Applied Mathematics, University of Western Onta-

     rio, London, Canada N6A 5B7 rcorless@julian.uwo.ca or rmc@pineapple.apmaths.uwo.ca

Publicity Chair: Gerhard Schneider, GWDG, D-37077 Göttingen, Germany. gschnei2@gwdg.de Treasurer: Alfred Widiger, FB Informatik, Universität Rostock, D-18051 Rostock, Germany.

     wid@informatik.uni-rostock.de

Local Arrangements Chair: Karl Hantzschmann, FB Informatik, Universität Rostock, D-18051 Rostock, Germa-

     ny. kh@informatik.uni-rostock.de

Exhibit Chair: Joachim Apel, Institute for Computer Science, University of Leipzig, Augustusplatz 10-11, D-04109

     Leipzig, Germany. apel@informatik.uni-leipzig.de

Proceedings Editor: Oliver Gloor, University of Tübingen WSI, Sand 13 D-72076 Tübingen, Germany.

     gloor@informatik.uni-tuebingen.de

Call for Posters: The Poster Sessions are an ideal venue for presenting recent research results or ongoing research projects that might not yet be complete, but whose preliminary results are already interesting nonetheless. Posters describing computer algebra systems and applications are also especially welcome. To encourage poster submissions of good quality, a best poster prize, consisting of a certificate, will be awarded. Posters will be judged based on content and presentation.

Instructions for authors: Authors should submit to the Poster Committee Chair an abstract of no more than 1000 words describing the contents of the poster. Plain LaTeX or TeX is preferred, but plain text is also acceptable. E-mail submissions are strongly encouraged and should be sent to glabahn@daisy.uwaterloo.ca Regular mail submissions should be sent via air mail to George Labahn, Symbolic Computation Group, Department of Computer Science, University of Waterloo, Waterloo, Ontario, CANADA N2L 3G1.

Please allow at least one week for air mail delivery.

Important Dates:

16 June 1998: Poster abstracts must be received
1 July 1998:  Notification of acceptance

Authors are encouraged to read the article How to Prepare a Poster http://www.siam.org/siamnews/general/poster.htm

Poster Sessions Committee: Poster committee members will be announced on the ISSAC’98 www-site.

The Vice-Chair of the Committee is Mohamed O. Rayes, Texas Instruments, mrayes@ti.com .

Further information about the symposium is available at http://wwwteo.informatik.uni-rostock.de/ISSAC98

11. ICM98 – International Conference in Mathematics Berlin, 18.8. – 28.8.98

The International Congresses of Mathematicians, taking place roughly every four years since 1897, belong to the most important mathematical events in the world. One distinguishing feature, among others, is the award of the Fields Medals and the Nevanlinna Prize (the ”mathematical Nobel Prizes”) during the opening ceremony.

The ICM’98 will take place in Berlin, Germany, from August 18 to 27, 1998. In addition to the scientific program (with plenary and invited speakers chosen by the IMU-appointed ICM’98 Program Committee) a ”Section of Special Activities” is planned. One of these activities will be a session on mathematical software, to be held on two afternoons during the congress. The focus of this session will be the presentation of a broad spectrum of mathematical software systems ranging from general purpose systems to specialized systems, e.g., systems from numerical analysis, computer algebra, optimization, mathematical visualization, or mathematical education. The presentations should include typical applications.

This session is planned to attract a broad audience: ICM attendees, students, teachers, etc., with a special interest in mathematical software. The session will take place at the conference site of the ICM’98 congress at the Technische Universitaet Berlin.

Mathematical Software Session

A program committee for this particular session has been appointed. It will be chaired by Johannes Grabmeier, IBM Germany (speaker of the special interest group for computer algebra of DMV (German Mathematical Society), GAMM and GI)

Organization: Winfried Neun, Konrad-Zuse-Zentrum fuer Informationstechnik, Berlin, Germany, Email: neun@zib.de .

Call for Presentations: The systems to be presented should meet the highest standards with respect to mathematical content. Mathematical originality, new solutions, or uncommon fields of application will be highly appreciated. The technical quality in design and implementation is also an important issue.

Submissions for the Session on Mathematical Software are encouraged from all fields of mathematics where software systems are used. Systems which are available free of charge (e.g., public domain) are especially invited and clearly preferred when the presentations will be chosen.

There will be a software exhibition and a book fair in connection with the ICM’98 congress, too. This may be more suitable for the demands of vendors of commercial software systems. Please contact the chairman of the local arrangements committee, Prof. Rolf H. Moehring, email: moehring@math.tu-berlin.de for details about the exhibition.

Talks in which various commercial packages are compared from an independent viewpoint, pointing out particular strengths and weaknesses of the systems, are also sought.

The program committee, a group of internationally renowned mathematicians and experts on mathematical software, will evaluate the entries and select a number of contributions according to quality and thematic balance. To facilitate adequate judgement, a submission should include material (either in paper form or an electronically readable format, e.g., an URL) which explains to the committee the mathematical background of the systems, the fields of application and the software design and techniques.

Submissions should be sent, preferably by electronic mail, to: ICM’98 – Session on Math. Software, c/o W. Neun, Konrad-Zuse-Zentrum (ZIB), Takustr. 7, D-14195 Berlin, Germany, Email: neun@zib.de and must be received by March 1, 1998 (Deadline for Submission).

Submissions that arrive after this deadline will not be considered. Some guidelines that will help the program committee to review the submissions are:

For a first glance a URL is usually very helpful.

For each system it should be very clear where information about the mathematical content can be found. This is usually not trivial if the submission consists, say, of uncommented pictures.

The special features and the targetted user community should be identified.

The availability of the software and the terms and conditions for distribution should be easily accessible.

The scheduled length of the presentations including discussion is 30 minutes. This allows the organizers to put approx. 12 lectures into the time available for the session.

Financial support for presentations is not available. Presenters are required to register for ICM’98.

Upon Acceptance: Contributors will be notified of the acceptance or rejection of their submission by the program committee. Based on this selection, the organizing committee will arrange a timetable in cooperation with the contributors of the presentations.

Demands for special equipment needed for presentations can be discussed at this time, but the resources will be limited. Therefore, it is not advisable to rely on any special hardware and software support from the session organizers.

It is the contributor’s responsibility to secure any necessary permissions and licenses for any material contained in his presentation or handouts. The organizers of ICM’98 would appreciate it if the commercial attitude of the system providers were modest.

Deadlines: Submission of Presentations March 1, 1998 Notification of Acceptance April 1, 1998

Other Topics of th ICM98: 1. Logic , 2. Algebra: Finite and infinite groups. Rings and algebras. Representations of finite dimensional algebras. Algebraic K-theory. Category theory and homological algebra. Computational algebra. Geometric methods in group theory. 3. Number Theory and Algebraic Arithmetic , 4. Algebraic Geometry (joint piece with 11) , 5. Differential Geometry and Global Analysis , 6. Topology , 7. Lie Groups and Lie Algebras ,

8. Analysis , 9. Ordinary Differential Equations and Dynamical Systems , 10. Partial Differential Equations (includes non-linear functional analysis) , 11. Mathematical Physics (joint piece with 4) , 12. Probability and Statistics , 13. Combinatorics , 14. Mathematical Aspects of Computer Science (joint with IUCSI) , 15. Numerical Analysis and Scientific Computing , 16. Applications , 17. Control Theory and Optimization (joint with Mathematical Programming Society) , 18. Teaching and Popularization of Mathematics , 19. History of Mathematics.

Further information on the conference can be found on World Wide Web at the URL: http://elib.zib.de/ICM98

12. Algebraic Number Theory and Diophantine Analysis Graz, Österreich, 30.8. – 5.9.98

This international conference is organized by

Franz Halter-Koch (University of Graz), Institut für Mathematik, Karl-Franzens Universität Graz, Heinrichstrasse 36, A-8010 Graz, Austria.

Robert F. Tichy (Graz Technical University), Institut für Mathematik A, Technische Universität Graz, Steyrergasse 30, A-8010 Graz, Austria.

The topics of the conference include algebraic number theory, diophantine equations, transcendence, uniform distribution as well as computational and analytic aspects. There will be one hour survey lectures as well as 20 minutes contributed talks (open for everybody) and a special session on diophantine equations.

The following mathematicians have already agreed to attend the conference:

Jozsef Beck (Rutgers University) Jörg Brüdern (University of Stuttgart) Jan-Hendrik Evertse (University of Leiden) Ernst-Ulrich Gekeler (University of Saarbrücken) Kalman Gyoery (University of Debrecen) Stephane Louboutin (University of Caen) Wladislaw Narkiewicz (University of Wroclaw) Andrew M. Odlyzko (Bell Labs), tentatively Attila Pethoe (University of Debrecen) Florian Pop (University of Bonn) Andrzej Schinzel (Polish Academy of Sciences), tentatively Wolfgang M. Schmidt (University of Colorado), tentatively Rene Schoof (University of Rome) Martin Taylor (University of Manchester) Robert Tijdeman (University of Leiden) Michel Waldschmidt (Institut Poincaré, Paris)

Graz is the capital of Styria, a southern province of Austria. The meeting will take place at the University and the Technical University, which are both in walking distance from the city center. Graz can be reached either by train or by plane. There are flights from Vienna, Zurich and Frankfurt.

There will be a possibility for moderately priced housing in a dormitory. Of course, it is possible to choose any hotel in Graz via the tourist office. The conference fee is approximately 1000 ATS (ca. 150DM, ca. 90US$ , ca. 450FF). In special cases a reduction of the conference fee may be possible. For more details we refer to the second announcement in approximately one year.

Everybody interested in the second announcement should contact the e-mail address:nt98@weyl.math.tu-graz.ac.at

13. Representation Theory of Algebras Bielefeld, 31.8. – 5.9.98

An International Conference devoted to Representation Theory of Algebras will be held August 31 - September 5, 1998 at the University of Bielefeld (starting on August 31 at 9:00 am and ending on September 5 at 2:00 pm).

The scientific organization lies in the hands of Dieter Happel (Chemnitz), Helmut Lenzing (Paderborn), Claus Michael Ringel (Bielefeld) and Klaus W. Roggenkamp (Stuttgart).

The local organizing committee consists of Thomas Brüstle, Peter Dräxler, Steffen König, Henning Krause, Rainer Nörenberg, Claus Michael Ringel and Jan Schröer.

Program: There will be 18 one-hour lectures on important recent developments in the subject and on decisive relations to other parts of mathematics. The following mathematicians have accepted our invitation for such a lecture:

Hans Joachim Baues (Bonn, Germany), Raymundo Bautista (UNAM, Mexico), Jon Carlson (Athens, USA), William Crawley-Boevey (Leeds, England), Gordon James (London, England), Mikhail Kapranov (Boston, USA), Olivier Mathieu (Strasbourg, France), Sergei Ovsienko (Kiev, Ukraine), Idun Reiten (Trondheim, Norway), Andrzej Skowronski (Torun, Poland), Katsuhiro Uno (Osaka University, Japan), Michel Van den Bergh (Hasselt, Belgium), Jie Xiao (Beijing, China).

The remaining one-hour lectures and additional twenty-minutes lectures (in parallel sessions) will be arranged on the basis of abstracts provided by participants.

On Tuesday evening (September 1) there will be a presentation of computer algebra packages which handle problems in representation theory, organized by Peter Dräxler (Bielefeld).

The final program will be available July 31, 1998 in the internet.

Abstracts The deadline for the submission of an abstract is June 30, 1998. Such an abstract should be no longer than one page and should clearly state the result to be presented. The abstracts obtained in time will be distributed to all participants.

Registration Fee There will be a registration fee of DM 200,- (for students: DM 150,-). It will cover the scientific material to be distributed as well as lunches (Monday - Friday), an excursion on Wednesday afternoon (September 2) and a joint dinner on Thursday evening (September 3). The registration fee will be reduced for those paying before June 30, 1998 to DM 150.- (and DM 100,- for students); the money has to be sent to the following bank account: Ringel, Sparkasse Bielefeld, BLZ 480 501 61, Account Number 43 00 75 41.

Accomodation: Several hotels in Bielefeld provide special rates for the participants of the conference, see the enclosed list. Every participant is asked to contact directly one of these hotels, or else to make an arrangement with: Tourist Information, Am Bahnhof, D-33602 Bielefeld, Germany. Tel: (+49) 521 178844, Fax: (+49) 521 178811.

The Week Before: The International Congress of Mathematicians (ICM 98) will be held at Berlin August 18 - 27, 1998. It is easy to go from Berlin to Bielefeld: one may use a train, it runs every hour and takes 210 minutes.

The Week After: This meeting will be followed by a Euroconference devoted to ”Infinite Length Modules”, September 7 - 11, 1998, also hosted by the University of Bielefeld (see the announcement below). Special funds of the European Union (TMR) are available for allowing young researchers from member states of the European Union to participate at the Euroconference. For those who register for both conferences (The International Conference on ”Representation Theory of Algebras” and the Euroconference on ”Infinite Length Modules”) we plan an excursion on Sunday, September 6.

Further information: http://www.mathematik.uni-bielefeld.de/~sek/

or you may contact: Ms. H. Scharsche, Fakultät für Mathematik, Universität, POBox 100 131, D-33501 Bielefeld, Germany, Fax: (+49) 521 106-4743, E-mail: scharsch@mathematik.uni-bielefeld.de .

14. Euro-Par’98 Southampton, 1.9. –4.9.98

Scope of the Conference: EURO-PAR is the annual European conference in Parallel Processing. The conference will consist of a panel of highly focused workshops on all aspects of parallel processing, from theory to practice, from academy to industry.

Subjects of the Workshop on Symbolic Computation: Papers are invited on all topics related to parallel symbolic computation. Symbolic computation is defined here in its broadest sense, comprising all programming and computation paradigms that deal with symbolic data, or that deal with exact numeric data rather than floatingpoint values. This includes functional programming, logic programming, constraint programming, computer algebra, theorem proving, AI paradigms, etc. Topics of interest comprise parallel execution models, parallel algorithms, parallel compilation and run-time techniques as applied to the specific domain of symbolic computation, including parallelizing recursion, speculative parallelism, manipulation of complex parallel data structures, parallel search, parallel language constructs, and static, dynamic or hybrid resource control (heap management, load management, data placement, granularity, etc.). Papers reporting on significant applications are particularly welcome.

Conference Organization Chair: Tony Hey Committee Members: Jeff Reeve, David Pritchard, Hugh Glaser. Steering Committee: Ron Perrott (Queen’s U. Belfast, Chair), Emilio Zapata (U. Malaga, Vice-Chair), Luc Bouge (ENS Lyon), Paul Feautrier (U. Versailles), Luco Grandinetti (U. Calabria), Seif Haridi (SICS), Peter Kacsuc (KFKI), Christian Lengauer (U. Passau), Karl Dieter Reinartz (U. Erlangen), Paul Spirakis (CTI), Marian Vajtersic (Slovak Academy), Richard Wait (MidSweden U.), Makoto Amamiya (Kyushu U.), Agnes Bradier (EC), Ian Foster (Argonne).

Workshop Organization General Chair: Wolfgang Küchlin, Universität Tübingen Local Chair: Ali Abdallah, University of Reading, England. Vice Chairs: to be announced.

Further information: http://www.europar98.ecs.soton.ac.uk

15. Infinite Length Modules Bielefeld, 7.9. – 11.9.98

A Euroconference Infinite Length Modules will be held September 7 - 11, 1998, at the University of Bielefeld (starting on September 7 at 9:00 am and ending on September 11 at 17:00 pm). The scientific organization lies in the hands of Ken Brown (Edinburgh), Paul M. Cohn (London), Idun Reiten (Trondheim) and Claus Michael Ringel (Bielefeld). The local organizing committee consists of Thomas Brüstle, Peter Dräxler, Steffen König, Henning Krause, Rainer Nörenberg, Claus Michael Ringel and Jan Schröer. Program The aim of the conference is to provide a survey of methods and results concerning modules of infinite length (in comparison to the well-known properties of finite length modules). A detailed program will be available by the end of February 1998. Registration Fee There will be a registration fee of DM 140,-. The registration fee will be reduced for those paying before June 30, 1998 to DM 100.-; the money has to be sent to the following bank account: Ringel, Sparkasse Bielefeld, BLZ 480 501 61, Account Number 43 00 75 41. Accomodation Several hotels in Bielefeld provide special rates for the participants of the conference, see the enclosed list. Every participant is asked to contact directly one of these hotels, or else to make an arrangement with: Tourist Information, Am Bahnhof, D-33602 Bielefeld, Germany. Tel: (+49) 521 178844, Fax: (+49) 521 178811. Financial Support The Euroconference is financially supported by the TMR-program of the European Union. Special funds are available for allowing young researchers from member states of the European Union to participate (see the registration form).

Further information: http://www.mathematik.uni-bielefeld.de/~sek/

or you may contact: Ms. H. Scharsche, Fakultät für Mathematik, Universität, POBox 100 131, D-33501 Bielefeld, Germany, Fax: (+49) 521 106-4743, E-mail: scharsch@mathematik.uni-bielefeld.de .

16. AISC’98 – The Fourth International Conference Artificial Intelligence and Symbolic Computation Plattsburgh, New York, USA 16.9. – 18.9.98

About this Conference Series: Conferences in this series are held every two years; the previous three took place in Karlsruhe (Germany), Oxford (United Kingdom) and Steyr (Austria). This time the conference will be held in USA and will be sponsored by the Plattsburgh State University of New York. Papers submitted to the conference undergo a standard review process. Previous proceedings appeared in the Springer Verlag LNCS Series, volumes 737, 958 and 1138. This year’s proceeedings are also expected to be published by Springer Verlag and available at the time of the conference. Later, expanded versions of best papers will appear after a review in a special volume of Fundamenta Informaticae. The aim of the conference is to provide a forum for the exchange of novel ideas and the presentation of latest solutions. Another goal is to foster personal contacts among researchers from diverse fields surrounding AI and symbolic computation. The conference is concerned with all aspects of research, including theory, implementations and applications. Topics: AI in Symbolic Mathematical Computing, Computer Algebra Systems and Automated Theorem Provers, Integration of Logical Reasoning and Computer Algebra, Engineering, Industrial and Operations Research Applications, Foundations and Complexity of Symbolic Computation, Mathematical Modeling of Multi-Agent Systems, Programming Languages for Symbolic Computation, Symbolic Computation for Expert Systems and Machine Learning, Implementations of Symbolic Computation Systems. Papers involving other topics but with significant links to the topics above will also be appropriate. Invited Speakers: Luc DeRaedt (Katholieke Universiteit Leuven, Belgium): Inductive logic programming. Melvin Fitting (City University of New York, USA): Bertrand Russell, Herbrand’s theorem, and the assignment statement. Richmond Thomason (University of Pittsburgh, USA): Repesenting and reasoning with context. Deadlines: Deadline for submissions: March 25, 1998, Information about acceptance of papers: May 1, 1998, Final version due: June 1, 1998. Submission Requirements: Authors are invited to submit papers up to 12 pages long, in English, conforming to the editorial requirements of the Springer Verlag LNCS series, formatted in LaTeX. (See http://www.springer-ny.com/samples/tools.html or contact organizers to receive LaTeX lncs.sty file and instructions). The title page should contain: title, author(s) with affiliation(s), e-mail address(es), specification of topic(s) from the list above to which the paper is related, listing of keywords, abstract. Results must be unpublished. Electronic submission is encouraged. Send both a postscript file and the LaTeX source code to aisc98@splava.cc.plattsburgh.edu (If electronic submission is not possible at this time, please send four hard copies to the address below.) Steering Committee: Jacques Calmet (Universitaet Karlsruhe, Germany), John Campbell (University College London, Great Britain), Jochen Pfalzgraf (Universitaet Salzburg, Austria), Jan Plaza (SUNY/Plattsburgh, USA), Conference Chairman. Program Committee: Luigia C. Aiello (Roma, Italy), Francois Arlabosse (Framatome, France), Bruno Buchberger (Linz, Austria), Gregory Butler (Montreal, Canada), James Cunningham (London, Great Britain), John Debenham (Sydney, Australia), Ruediger Dillmann (Karlsruhe, Germany), Fausto Guiunchiglia (Trento, Italy), Stan Klasa (Montreal, Canada), Alfonso Miola (Roma, Italy), Lin Padgham (Melbourne, Australia), Zbigniew W. Ras (Charlotte, USA), Klaus U. Shultz (Munich, Germany), Joerg H. Siekmann (Saarbruecken, Germany), Andrzej Skowron (Warsaw, Poland), Stanly Steinberg (Albuquerque, USA), Karel Stokkermans (Linz, Austria), Carolyn Talcott (Stanford, USA), Peder Thusgaard Ruhoff (Odense, Denmark), Dongming Wang (Grenoble, France).

Addresses: Up-to-date information is always available on the web page http://www.plattsburgh.edu/aisc98/

Submissions and questions can be directed via e-mail to: aisc98@splava.cc.plattsburgh.edu, Phone: (518) 564 2788, Fax: (518) 564 3010,

Mail can be addressed to: AISC’98,, c/o Jan Plaza, Computer Science Department, Plattsburgh State University of New York, Plattsburgh, NY 12901, USA.

17. Cooperating Systems for Symbolic Computation Potsdam, 23.9. – 25.9.98

Research in Symbolic Computation and its fields Computer Algebra, Deduction and Model Checking has led to the development of quite a few systems that -taken together- produced remarkable results. Although some of these systems are capable to deal with many problems, even if we take only a subfield there is not one system that can deal with all the problems users are interested in. This is the reason why there are so many specialized systems that solve certain tasks very efficiently.

So, at the moment there are some systems that have a general purpose intend (at least for one of the fields of Symbolic Computation) and that were very costly in their development. On the other hand, there are many special purpose systems that solve certain very difficult tasks with specialized methods. Naturally, users are interested in systems that combine all the best of existing technologies.

One solution to this problem is to let the existing systems cooperate. In the last few years this solution has been the subject of the research of several groups in symbolic computation. In order to be useful, systems for symbolic computation must be also capable to cooperate with a wide range of application systems from other fields. These include text editors as well as data bases, verification systems and many others.

The basis for the cooperation of systems is the cooperation of the people that design and implement them. And the basis for cooperation is the exchange of information. Therefore it is the time to bring together the researchers interested in Cooperative Symbolic Computation Systems in order to exchange ideas and experience and to provide the basis for extended cooperations of both people and systems.

Topics: The topics of interest of the workshop include but are not limited to: cooperation concepts for automated and interactive deduction, computer algebra, model checking, and combinations of these fields, cooperation of several symbolic computation systems and human beings, cooperation of systems for symbolic computation with systems from other fields, and interchange formats for symbolic computation

We especially encourage submissions describing systems or case studies to these topics.

Important Dates: Papers Due: April 1, 1998, Author Notificaton: After May 22, 1998, Camera Ready Copy Due: June 22, 1998, Workshop: September 23-25, 1998.

Submission Information: The submitted paper should not exceed 15 pages. We strongly encourage to use the standard LaTeX article style. The primary means of submission will be electronic by mailing an uuencoded compressed (or gziped) PostScript file to one of the program co-chairs. Submissions received after March 15 will not be considered.

Proceedings: Accepted contributions will be published in a technical report of the Computer Science Department of the University Kaiserslautern. The report will be available at the workshop.

Program Committee:

General Chair: Bruno Buchberger, RISC Linz, Austria.

Program Committee Co-Chairs: Jacques Calmet, Faculty for Computer Science, Karlsruhe University, Germany, calmet@ira.uka.de . Ingo Dahn Mathematical Institute, Humboldt University Berlin, Germany dahn@mathematik.hu-berlin.de

Program Committee Members: Jörg Denzinger, University Kaiserslautern, Germany, John Harrison, University of Cambridge, UK, Tudor Jebelean, RISC-Linz, Austria, Tobias Nipkow, TU Munich, Germany, Laurent Thery, INRIA, France, Andreas Weber, University of Tübingen, Germany.

Local Organization Chair: Martin Weese, Mathematical Institute, Potsdam University, Germany weese@rz.uni-potsdam.de

Actual Information:

See http://www.uni-kl.de/AG-AvenhausMadlener/cosysyco.html

18. Algorithmen und Zahlentheorie Dagstuhl, 12.10. – 30.10.98

Leiter: H. Niederreiter, Academy of Sciences, Vienna, Austria, A. M. Odlyzko, AT&T Florham Park, USA, M. E. Pohst, TU Berlin, Germany.

Berichte von Konferenzen

1. Algorithmische Ideal- und Invariantentheorie (DMV-Seminar) Oberwolfach, 12. – 18.10.1997

Das von G.-M. Greuel und B. H. Matzat geleitete DMV-Seminar fand mit 26 Teilnehmern statt. Darin wurde eine Einführung in die algorithmische Idealtheorie gegeben mit Anwendungen im Bereich der Invarianten- und Singularitätentheorie. Dazu wurden Übungen veranstaltet und Computerpraktika hauptsächlich mit SINGULAR und MAGMA durchgeführt. Die einzelnen Vorträge (in zeitlicher Reihenfolge) waren:

G.-M. Greuel (Kaiserslautern), Einführung von Gröbner- und Standardbasen I, II G. Pfister (Kaiserslautern), Primärzerlegung und Normalisierung I, II B. H. Matzat (Heidelberg), Invariantenringe endlicher Gruppen I, II G. Kemper (Heidelberg), Berechnung von Invariantenringen B. H. Matzat, Invariantenringe linear reduktiver Gruppen I, II G. Kemper, Invarianten endlicher Spiegelungsgruppen G.-M. Greuel, Singularitäten B. H. Matzat, Generische Polynome und das Noethersche Problem G. Kemper, Tiefe von Invariantenringen

Eine Ausarbeitung des DMV-Seminars in Buchform (für den Birkhäuser-Verlag) ist in Vorbereitung.

B. H. Matzat (Heidelberg)

2. Algorithmic Aspects of Drinfeld Modules Dagstuhl, 1.12. – 6.12.97

Dieser von E.-U. Gekeler und B. H. Matzat geleitete (letzte) Workshop im Rahmen des DFG-Schwerpunktprogramms Algorithmische Zahlentheorie und Algebra wurde von 34 Teilnehmern besucht. Es wurden die folgenden Vorträge (in zeitlicher Reihenfolge) gehalten:

H. G. Rück (Essen), L-series of Heegner points of Drinfeld modular curves U. Tipp (Saarbrücken), Green functions for Drinfeld modular curves A. Bruns (Essen), Explicit equations of Drinfeld modular curves M. Traulsen (Essen), Modular symbols for function fields A. Schweizer (Montreal), Elliptic curves over F_{q}(T) with conductor of degree four G. Cornelissen (MPI Bonn), Zeroes of Eisenstein series I N. Rebinguet (Toulouse), Zeroes of Eisenstein series II E.-U. Gekeler (Saarbrücken), Zeroes of Eisenstein series III Cho Xing (Wien), Construction of algebraic curves over finite fields with many rational points B. Wack (Saarbrücken), Zeta functions of elliptic curves over F_{2}(T) O. Scheja (Saarbrücken), Zeta functions of modular graphs S. Abhyankar (Purdue), Galois theory of Moore-Carlitz Drinfeld modules B. Anglès (Toulouse), On the analogue of a problem of Mirimanoff in characteristic p. D. Thakur (Univ. of Arizona), Results and computational projects in the arithmetic of Drinfeld modules L.-Ch. Hsia (Nat. Cent. Univ. Taiwan), On characteristic polynomials of geometric Frobenius associated to Drinfeld modules I. Potemine (KIAS Seoul), Drinfeld-Anderson motives B. Lopez (Madrid), On the expansion of a product formula related with the Drinfeld discriminant function F. Recher (Lille), Transcendence of the modular invariant in finite characteristic R. Pink (Mannheim), The Galois action on the torsion points of a Drinfeld module J. Yu (Acad. Sinica), On class number relations over function fields D. Hayes (Amherst), j-invariants of CM-Drinfeld modules D. Dorman (Middlebury), Explicit factorization formulae for singular moduli of Drinfeld modules S. Bae (KAIST), On the coefficients of Drinfeld modular equations

Der Workshop wurde mit einer Problem Session abgeschlossen.

B. H. Matzat (Heidelberg)

3. RISC Linz: Intensive Course on Gröbner Bases Linz, Österreich, 7.1. – 30.1.98

The Intensive Course on Gröbner Bases, organized by Prof. Bruno Buchberger and Prof. Franz Winkler, took place in Hagenberg, the home of RISC-Linz, from January 12 through 30. The course was primarily designed as a comprehensive introduction for ph.d. students and young postdocs. The 28 participants came from Austria, Belarus, Czechia, Germany, Hungary, Iran, Italy, Japan, Moldavia, Norway, Romania, Russia, Spain, Sweden, Ukraine, and USA.

List of lecturers and subjects:

B. Buchberger (Linz): Introduction to the theory and construction of Gröbner bases, elementary applications such as ideal membership, computation in residue class rings, solution of algebraic equations, linear syzygies, Hilbert functions, non-linear syzygies, and inverse mappings. F. Winkler (Linz): Introduction to general rewriting, reductions relations, term rewriting, geometric applications of Gröbner bases, such as implicitization and parametrization of algebraic varieties, computation of dimension of a variety. F. Schwarz (Bonn): Janet bases as an algebraic tool for investigating partial differential equations, applications of Janet bases for vector fields of Lie groups, for 2nd order ordinary differential equations. L. Robbiano (Genova): Introduction to the theory of Hilbert functions and the computation of Hilbert functions, how to compute Gröbner bases with the aid of Hilbert functions, using Hilbert functions for computing free resolutions. W.W. Adams (College Park, MD): Gröbner bases over rings, primality testing using Gröbner bases, primary decomposition of ideals by Gröbner bases.

B. Buchberger and F. Winkler (Linz)

4. International Conference on Gröbner Bases (“33 Years of Gröbner Bases”)

Linz, Österreich, 2.2. – 4.2.98

The conference “33 Years of Gröbner Bases” was held in Hagenberg near Linz, the home of RISC-Linz, from February 2 through 4. Prof. Bruno Buchberger and Prof. Franz Winkler (both RISC-Linz) were the chairmen of the conference. The 94 participants came from 14 countries.

The conference consisted of one day of tutorial lectures on various applications of the theory of Gröbner bases, and two days of presentations of original research. The proceedings are published as: B.Buchberger and F.Winkler (eds.), Gröbner Bases and Applications, London Math. Soc. LNS 251, Cambridge Univ. Press, 1998.

Titles of the tutorial lectures and research talks in alphabetic order of the authors follow.

Tutorial lectures: B. Buchberger (Linz): Introduction to Gröbner Bases. F. Chyzak (Rocquencourt): Gröbner Bases, Symbolic Summation and Symbolic Integration. W. Decker and T. de Jong (Saarbrücken): Gröbner Bases and Invariant Theory. G.-M. Greuel and G. Pfister (Kaiserslautern): Gröbner Bases and Algebraic Geometry. S. Hoşten and R. Thomas (Fairfax, VA and College Station, TX): Gröbner Bases and Integer Programming. H.M. Möller (Dortmund): Gröbner Bases and Numerical Analysis. L. Robbiano (Genova): Gröbner Bases and Statistics. D. Wang (Grenoble): Gröbner Bases Applied to Geometric Theorem Proving and Discovering.

Original research talks: B. Amrhein and O. Gloor (Bern and Tübingen): The Fractal Walk. M.A. Borges and M. Borges (Santiago de Cuba): Gröbner Bases Property on Elimination Ideal in the Noncommutative Case. M.-J. Gonzalez-Lopez and L. Gonzalez-Vega (Santander): Newton Identities in the Multivariate Case: Pham Systems. M. Insa and F. Pauer (Innsbruck): Gröbner Bases in Rings of Differential Operators. J.B. Little (Worchester, MA): Canonical Curves and the Petri Scheme. H. Lombardi and H. Perdry (Besançon): The Buchberger Algorithm as a Tool for Ideal Theory of Polynomial Rings in Constructive Mathematics. K. Madlener and B. Reinert (Kaiserslautern): Gröbner Bases in Non-Commutative Reduction Rings. J.L. Miller (Bowling Green, KY): Effective Algorithms for Intrinsically Computing SAGBI-Gröbner Bases in a Polynomial Ring over a Field. Müller-Quade, R. Steinwandt and T. Beth (Karlsruhe): An Application of Gröbner Bases to the Decomposition of Rational Mappings. P. Nordbeck (Lund): On some Basic Applications of Gröbner Bases in Non-commutative Polynomial Rings. L. Robbiano and M.P. Rogantin (Genova): Full Factorial Designs and Distracted Fractions. T. Sauer (Erlangen): Polynomial Interpolation of Minimal Degree and Gröbner Bases. J. Schicho (Linz): Inversion of Birational Maps with Gröbner Bases. J. Snellman (Stockholm): Reverse Lexicographic Initial Ideals of Generic Ideals are Finitely Generated. Q.-N. Tran (Linz): Parallel Computation and Gröbner Bases: An Application of Converting Bases with the Gröbner Walk.

The following invited talks were also given at the conference: W.W. Adams (College Park, MD): Some Applications of Computational Algebra to Partial Differential Equations. B. Buchberger (Linz): 33 Years of Gröbner Bases: The Early Years. F. Schwarz (Bonn): Janet Bases for Symmetry Groups.

F. Winkler (Linz)

5. Computeralgebra in Forschung und Lehre, Workshop Künzelsau, 5.3. – 6.3.98

Der Workshop Computeralgebra in Forschung und Lehre in Künzelsau wurde veranstaltet von der FG Computeralgebra der DMV, GI und GAMM sowie der Fachhochschule Heilbronn. Die Örtliche Tagungsleitung hatte Wilhelm Werner (werner@fh-heilbronn.de). Mehr als 50 Teilnehmer aus ganz Deutschland beteiligten sich an dem zweitägigen Workshop, in dem zahlreiche Einsatzmöglichkeiten von Computeralgebra vorgestellt und lebhaft diskutiert wurden. Mehrere Verlage und Distributoren von einschlägiger Software nahmen die Gelegenheit wahr, um ihre aktuellen Produkte vorzustellen; im Rahmenprogramm wurde die Ausstellung Italienische Transavantguardia im Museum Würth besucht. Das Vortragsprogramm: Michael Möller (Dortmund): Das Lösen von algebraischen Gleichungssystemen, Harald Gross (Ulm): Exakte und schnelle Rechenverfahren für Matrizen, Karl Buehler (Offenburg): Anwendungen der Computeralgebra in Thermo- und Fluiddynamik, Karl-Heinz Zimmermann (Hamburg-Harburg): Erzeugung guter Codes, Laurent Bernardin (Zürich): Computeralgebra, Maple und die Zukunft, Thomas Westermann (Karlsruhe): Computeralgebrasysteme in der Mathematikvorlesung, Michael Komma (Reutlingen): Mathematikunterricht und neue Technologien, Hans-Dieter Janetzko (Konstanz): Direkte Berechnung der Hertzschen Pressung mit MATHEMATICA, Edgar M.E.Wermuth (Nürnberg): Invariante Mannigfaltigkeiten mit Maple, Wolfram Koepf (Leipzig): Potenzreihenentwicklungen und Summation in der Computeralgebra, Burkhard Alpers (Aalen): Flexible elektronische VDI-Richtlinien am Beispiel von Kurvengetrieben, Günter Gramlich (Kaiserslautern): Scientific Computing in der Strahlentherapie, Hans-Gert Gräbe (Leipzig): Einführung in die Computeralgebra - eine Lehrveranstaltung zwischen Mathematik und Informatik, Axel Schenk (Isny): Computerbasierte Mathematik-Möglichkeiten und Probleme am Beispiel von MAT-LAB Im Jahr 2000 soll die Veranstaltungsreihe mit einem Workshop an der Fachhochschule Konstanz fortgesetzt werden; Interessenten wenden sich bitte an heinrich@fh-konstanz.de.

W. Werner (Künzelsau)

Themen und Anwendungen der Computeralgebra

Audiosignalverarbeitung

Michael Clausen, Frank Kurth Institut für Informatik V, Rheinische Friedrich-Wilhelms-Universität Bonn

                 clausen@cs.uni-bonn.de, kurth@informatik.uni-bonn.de

Die Audiosignalverarbeitung (ASV) entstand als eines der ersten Teilgebiete der Nachrichtentechnik. Mittlerweile ist aus dieser traditionellen Ingenieurwissenschaft ein interdisziplinäres Forschungsgebiet mit starken mathematischen, informatischen sowie wahrnehmungspsychologischen Komponenten geworden. Die Entwicklungen der letzten Jahre waren durch den Vormarsch des Internets und der Multimediakommunikation geprägt. Anwendungen wie Audio-on-demand, Online-Kompression und echtzeitfähige Kommunikationsdienste im Internet führten zu zahlreichen offenen Fragestellungen. Die Computeralgebra ist, besonders vom systematisch-strukturellen Standpunkt betrachtet, ein wichtiges Hilfsmittel, das bereits zur Lösung vieler Fragestellungen auf dem Gebiet der ASV beitragen konnte. Im folgenden soll die Rolle der Computeralgebra in der ASV anhand einiger Beispiele skizziert werden.

Elementare Bausteine der ASV lassen sich geeignet durch algebraische Strukturen beschreiben. Dies sind insbesondere Polynome, (formale) Potenz- und Laurentreihen sowie Matrizen über bestimmten Körpern oder Ringen, mit deren Hilfe Signale, Filter und lineare Systeme beschrieben werden. Viele dieser linearen Systeme tragen ein hohes Maß an Struktur und ermöglichen somit den Entwurf effizienter Algorithmen. Beispielhaft seien hier Fourier-, Cosinus- und (im Zusammenhang mit Wavelettransformationen) Block-Toeplitz-Transformationen genannt.

Diskrete Signale entstehen zumeist aus analogen Signalen durch zeitlich äquidistante Abtastung und lassen sich durch Samplefolgen (x_{j})_{j∈ZZ} beschreiben. Hierbei wird meist ein bestimmter Folgenraum, wie der Hilbertraum ℓ^{2}(ZZ) der quadratsummierbaren Folgen, zugrundegelegt. Jedoch sind für Anwendungen im allgemeinen nur Folgen mit endlichem Träger relevant. Realisierungen innerhalb digitaler Schaltkreise oder Rechenanlagen erfordern außerdem eine Quantisierung der einzelnen Samples, d.h. eine Einschränkung auf eine endliche Anzahl möglicher Werte (siehe z.B.[11]). Eine besonders häufig genutzte Darstellungsform für diskrete Signale verwendet Verzögerungselemente (engl. Delays), um den zeitlichen Verlauf von Signalen zu beschreiben. Ein Verzögerungselement beschreibt hierbei einen Zeitschritt. Diese Art der Darstellung eignet sich insbesondere zur Simulation konkreter (Hardware-) Realisierungen digitaler Schaltungen. Formal werden die Verzögerungselemente meist durch die komplexe Variable z dargestellt. Folgen werden vermöge (x_{j})_{j∈ZZ} ↦ ∑_{j} x_{j}z^{−j} =: X(z) in die Verzögerungsdarstellung, ihre z-Transformierte, überführt. Hierbei werden Verzögerungselemente multiplikativ geschrieben, wobei der Faktor z^{k} einer Änderung von k Zeitschritten entspricht. Negative Exponenten stellen Samples bzw. Ereignisse der Vergangenheit, positive dagegen solche aus der Zukunft dar. Innerhalb einer konkreten Signalverarbeitung modelliert dies die Tatsache, daß nur momentan oder in der Vergangenheit aufgetretene (oder beim Empfänger angekommene) Samples für die Berechnungen benutzt werden können. Es können also letztlich nur Systeme verwendet werden, die sich aus negativen Verzögerungselementen aufbauen. Solche Systeme heißen kausal. Eine besondere Bedeutung kommt der Fourierreihendarstellung eines Signals, ω ↦ X(e^{iω}) für ω ∈ [0, 2π), zu. Hier bestimmt die Anzahl Samples pro Zeiteinheit, die sog. Abtastrate, implizit den von der Fouriertransformation dargestellten Frequenzbereich, die Bandbreite [7] des Signals. Auf diese Weise kann sowohl das Frequenzverhalten von Signalen als auch das von Filtern bestimmt oder beschrieben werden.

Zur Beschreibung von Transformationen auf den Folgenräumen weicht man ebenfalls häufig auf den Raum der z-Transformierten aus. Für ℓ^{2}(ZZ) wird dies durch den Satz von Riesz-Fischer über die Isomorphie von ℓ^{2}(ZZ) und L^{2}(T), den z-Transformierten auf dem Einheitskreis, gerechtfertigt. Ein einfach zu beschreibender Operator auf dem Folgenraum ist z.B. eine Translation um k Samples. Im Raum der z-Transformierten entspricht dies einer Multiplikation mit z^{k}. Allgemeiner läßt sich die Wirkungsweise eines linearen zeitinvarianten Filters auf dem Folgenraum durch eine Faltung beschreiben. Im z-Bereich ist dies ausgehend vom Signal (x_{j})_{j∈ZZ} und Filter (t_{j})_{j∈ZZ} eine Multiplikation Y (z) = T (z)X(z). Kombination mehrerer Faltungen, Translationen sowie einiger weiterer Operatoren liefert allgemeiner Filterbänke. (Aus algebraischer Sicht handelt es sich bei Filterbänken um spezielle Darstellungen geeigneter Köcher.) Das Ziel ist hier die Zerlegung eines Eingangssignals in mehrere Subbandsignale, die verschiedene Frequenzanteile des Signals enthalten. Wichtige Klassen von Filterbänken entstehen durch Faktorisierungen bzw. Zerlegungen von T (z) für den Fall |T (z)| = 1, |z| = 1. Beispiele sind Filterbänke der Form z^{−ℓ} = T (z) = ∑_{j=0}^{n} S_{j}(z)A_{j}(z), ℓ ≥ 0, wobei (a_{j,k})_{k∈ZZ} das j-te Analysefilter bzw. (s_{j,k})_{k∈ZZ} das j-te Synthesefilter ist. Da Y (z) = z^{−ℓ}X(z) in diesem Fall nur eine Verzögerung des Eingangssignals beschreibt, nennt man solche Filterbänke auch Filterbänke mit perfekter Rekonstruktion. Methoden zur Behandlung von Polynomen und Potenzreihen, wie Faktorisierung oder schnelle Multiplikation, finden somit in Theorie und Entwurf von Filterbänken und Filterbankalgorithmen einen breiten Anwendungsbereich. Im Falle von Multiratenfilterbänken, dies sind Filterbänke, die zusätzlich Operatoren zur Abtastratenänderung verwenden, gibt es weiterhin einen natürlichen Zusammenhang zwischen den möglichen Filterbankfaktorisierungen und der linearen Darstellung des größten gemeinsamen Teilers der verwendeten Filter [19]. Die Qualität eines Filters hängt meist von der Gestalt seiner z-Transformierten ab. Z.B. sollte deren Betrag für Frequenzen, die abgedämpft werden sollen, möglichst klein sein. Dies steht wiederum in engem Zusammenhang mit den (komplexen) Nullstellen der z-Transformierten. Eine typische Designforderung an Filterbänke wäre etwa perfekte Rekonstruktion unter Vorgabe bestimmter Nullstellen. Ein bekanntes Beispiel sind die Waveletfilter von Daubechies [10].

Die Verarbeitung digitaler Audiosignale in CD-Qualität (44.1 kHz Abtastrate PCM, 16 Bit lineare Quantisierung) bringt relativ große Datenraten von ca. 1.5 MBit/Sek. mit sich. Viele Multimedia- und Internetanwendungen erfordern daher sowohl hinreichende Datenkompression als auch geeignete Schutzmechanismen zur Vermeidung von Übertragungsfehlern oder Paketverlusten. Speziell zur Datenkompression sind in den letzten zehn Jahren zahlreiche Standards entwickelt worden. Zwei der bekanntesten Standards im Audio- und Videobereich sind MPEG-1 und MPEG-2 von der Moving Pictures Experts Group (MPEG) der International Organisation of Standardisation (ISO), siehe z.B. [16]. Im Audiobereich liefern die verwendeten Verfahren durch gezielte Kombination psychoakustischer Modelle und Computeralgebra-Methoden (schnelle Fourier- und Cosinustransformation, schnelle Multiratenfilterung) Kompressionsraten von bis zu 1 : 12 unter Beibehaltung von CD-Qualität. Es ist zu vermuten, daß sich diese Kompressionsraten durch den Einsatz besserer Filterbänke und verfeinerter psychoakustischer Modelle noch weiter steigern lassen.

Die FFT wird neben der Signalanalyse zu zahlreichen anderen Aufgaben eingesetzt. Ein Beispiel ist das Filterdesign. So wird in [8] eine Methode zum FFT-gestützten Entwurf von sog. Equiripplefiltern beschrieben. Da schnelle Fouriertransformationen bekanntlich schnelle Faltungsalgorithmen liefern, sind hierdurch effiziente Filteralgorithmen, z.B. für echtzeitfähige Implementierungen auf Signalprozessorbasis, zu gewinnen. Übersichten über Algorithmen zur schnellen Fouriertransformation geben [2, 3, 6, 9]. Viele Aufgabenstellungen, wie der Filterentwurf, erfordern zusätzlich zur Effizienz möglichst genaue numerische Berechnungen. Besonders bei der Verwendung von Festkommasignalprozessoren kommt es durch die Akkumulierung von Rundungsfehlern häufig zu nicht akzeptablen Rechenfehlern. Ein neuer FFT-Algorithmus [5], der auf schneller Approximation komplexer Zahlen durch Elemente des für n ≥ 7 in ℂ dichten Teilrings ZZ[e^{2πi/n}] und der Verwendung chinesischer Restetechniken basiert, umgeht dieses Problem.

Zahlreiche Softwarepakete können zur Unterstützung von ASV-Anwendungen verwendet werden. Eines der momentan bekanntesten Systeme ist das Paket MatLab zur numerischen Signalverarbeitung und Visualisierung. Eine Erweiterung dieses Systems, die sog. Symbolic Toolbox, enthält die Kernfunktionalität des Computeralgebrasystems Maple, so daß auch symbolische Berechnungen aus MatLab heraus möglich sind. Des weiteren gibt es zahlreiche spezialisierte Softwarepakete, die zum großen Teil frei verfügbar sind, und auf MatLab aufsetzen. Beispiele sind das Paket WaveLab [4] für die Behandlung von Wavelets oder die FBT-Tools [13, 14] für Multiratenfilterbänke. Weitere Programmbibliotheken, die Hilfsmittel aus der Computeralgebra verwenden, sind z.B. Filterdesign-Routinen. Hier sei nur der Entwurf von Multiratenfilterbänke mit quasi perfekter Rekonstruktion erwähnt, wie sie auch in den MPEG Audiocodern verwendet werden. Das Filterentwurfsproblem führt für solche Filter häufig auf quadratische Optimierungsprobleme mit quadratischen Nebenbedingungen [15].

Wir hoffen, mit diesem kurzen Artikel einen ersten Eindruck möglicher Beiträge der Computeralgebra zum Gebiet der ASV gegeben zu haben. Als einführende Literatur seien die Bücher [11, 12, 20] sowie einige der einschlägigen Zeitschriften [1, 17, 18] erwähnt.

Literatur

[1] Audio Engineering Society (AES). Journal of the AES.

[2] Thomas Beth. Verfahren der schnellen Fourier-Transformation. Number 61 in Leitfäden der angewandten Mathematik und Mechanik. Teubner, 1984.

[3] Richard E. Blahut. Fast Algorithms for Digital Signal Processing. Addison Wesley, 1987.

[4] J. Buckheit, S. Chen, D.L. Donoho, and I. Johnstone. About WaveLab, 1996.

    ftp://playfair.stanford.edu/∼wavelab.

[5] Joe Buhler, M. Amin Shokrollahi, and Volker Stemann. Fast and Precise Computations of Discrete Fourier Transforms using Cyclotomic Integers. STOC’97, El Paso, Texas, USA, pages 40–47, 1997.

[6] C.S. Burrus. Notes on the FFT. Rice University, Houston, Texas, 1995.

[7] P.L. Butzer, W. Splettstößer, and R.L. Stens. The sampling theorem and linear prediction in signal analysis. Jahresbericht der Deutschen Mathematiker-Vereinigung, 90:1–70, 1988.

[8] A. Enis Çetin, Ömer N. Gerek, and Yasemin Yardimci. Equiripple FIR Filter Design by the FFT Algorithm. IEEE Signal Processing Magazine, pages 60–64, 1997.

[9] Michael Clausen and Ulrich Baum. Fast Fourier Transforms. BI Wissenschaftsverlag, 1993.

[10] I. Daubechies. Orthonormal bases of compactly supported wavelets. Comm. Pure Applied Math., XLI(41):909–996, 1988.

[11] N.S. Jayant and Peter Noll. Digital Coding of Waveforms. Prentice Hall, 1984.

[12] M. Kahrs and K. Brandenburg, editors. Applications of Digital Signal Processing to Audio and Acoustics. Kluwer, 1998.

[13] Frank Kurth and Michael Clausen. Adaptive FBT and M-band wavelet packet algorithms in audio signal processing. Submitted to the IEEE Transactions on Signal Processing, 1997.

[14] Frank Kurth and Michael Clausen. M-band wavelet packets and filter bank trees as flexible tools in audio signal processing. In Proc. IEEE WASPAA 97, Mohonk, New Paltz, NY, 1997.

[15] Truong Q. Nguyen. Digital filter bank design quadratic-constrained formulation. IEEE Transactions on Signal Processing, 43:2103–2108, 1995.

[16] Peter Noll. MPEG Digital Audio Coding. IEEE Signal Processing Magazine, pages 59–81, 1997.

[17] The Institute of Electrical and Electronics Engineers (IEEE). Transactions on Signal Processing.

[18] The Institute of Electrical and Electronics Engineers (IEEE). Transactions on Speech and Audio Processing.

[19] M. Vetterli and C.Herley. Wavelets and filter banks: Theory and design. IEEE, Trans. on ASSP, 9:2207–2232, 1992.

[20] U. Zölzer. Digitale Audiosignalverarbeitung. Teubner, 1996.

Neues über Systeme und Hardware

Maple V Release 5.0

Mit Release 5.0 führt Waterloo Maple die Entwicklung eines offenen, leistungsfähigen mathematischen Werkzeugs fort. Diese offene Umgebung befähigt den Anwender von Maple V, interaktiv auf eine große Zahl kommerziell erhältlicher Softwarepakete zuzugreifen. Ingenieure aus allen technischen Bereichen in der Industrie steigern somit ihre Produktivität und verkürzen Entwicklungszeiten.

”Ingenieure benötigen heutzutage eine Softwareumgebung, die ein breites Spektrum von Aufgaben löst”, erklärt Dieter Hensler, Präsident und CEO von Waterloo Maple Inc. ”Die aktuellen technischen Aufgabenstellungen werden immer komplexer und der Bedarf nach immer größerer Genauigkeit steigt. Aus diesem Grund werden die Forderungen nach einem Softwaresystem vom Kaliber Maple V immer lauter. Der symbolische Ansatz von Maple V zur Lösung komplexer Aufgabenstellungen ermöglicht den mathematischen Entwurf von Modellen, noch bevor viel Zeit in die Entwicklung teurer Prototypen investiert werden muß. Mit Maple V Release 5.0 stellt sich Waterloo Maple einmal mehr den Herausforderungen stetig wachsender Bedürfnisse.”

Release 5.0 ist die Lösung, die im technischen Umfeld gebraucht wird. Maple V ist anerkannt für seine Offenheit, seine technische Leistungsfähigkeit und für seine Genauigkeit. Die neuen Funktionen in Release 5.0 erlauben dem Anwender einfachste Dateneingabe und schnellen Zugang zu seinen Lösungen in einer bedienerfreundlichen und leicht erlernbaren Umgebung. Zum Beispiel können mit Release 5.0 Excel-ähnliche Tabellen definiert werden - eine Technologie, die schon von Millionen von Anwendern erfolgreich genutzt wird. Ab sofort stehen in Release 5.0 auch kontextabhängige Menüs sowie Paletten zur Eingabe von Formeln, Matrizen oder Sonderzeichen zur Verfügung. Diese Merkmale machen den Umgang mit diesem intuitiven ”click & solve” Werkzeug bedeutend bequemer.

Viele der neuen Leistungsmerkmale werden von Ingenieuren in der Industrie angewandt, die eine integrierte, offene Umgebung suchen. Der MATLAB-Link als Beispiel ermöglicht den direkten Zugang zu der numerischen Leistungsfähigkeit von MATLAB aus Maple V heraus. Release 5.0 besitzt weiterhin eine Funktion für den Export von Maple V Arbeitsblättern nach HTML. Damit können sich Anwender schnell über laufende Projekte informieren, Forschungsmaterial austauschen und technische Dokumentationen über das Inter- oder Intranet veröffentlichen. Kombiniert mit der ”click & solve” Technologie erweitern diese Funktionen die Modellierungs-, Lösungs- und Präsentationsmöglichkeiten enorm. Die neuen Leistungsmerkmale von Maple V Release 5.0 umfassen unter anderem: MATLAB Link für maximale Flexibilität

Release 5.0 beinhaltet einen direkten Zugriff auf MATLAB aus der Maple V Umgebung heraus. Mit der Kombination numerischer Leistungsfähigkeit von MATLAB und dem mathematischen Potential von Maple V läßt sich jetzt ein noch größeres Spektrum von Aufgaben lösen. Tabellenkalkulation inklusive

Innerhalb von Maple V Arbeitsblättern lassen sich Tabellen einfügen, in denen algebraische Ausdrücke verarbeitet werden können. Diese Tabellen besitzen die Grundfunktionalität von gängigen Tabellenkalkulationsprogrammen, wie z.B. Excel. Export nach HTML

Mit Hilfe von Maple V Release 5 können Anwender mit wenigen Mausklicks HTML Seiten erstellen. Alle Ausdrücke in mathematischer Formelnotation als auch Grafiken sind hierin enthalten. Jetzt lassen sich mathematisch-technische Online-Dokumentationen für räumlich übergreifende Projekte oder Online-Präsentationen auf Knopfdruck erstellen. Wesentlich verbesserte Benutzerschnittstelle

Maple V gehört unbestritten zu den führenden mathematischen Softwarepaketen, insbesondere in Bezug auf einfache Bedienung und leichtes Erlernen. Das Arbeitsblatt-Konzept garantiert den effizientesten Weg für die Eingabe und Ausgabe mathematischer Informationen und bietet umfangreiche Dokumentationsmöglichkeiten. Einfachste Eingabe mit Paletten

Maple V Release 5.0 enthält eine Reihe von Paletten, die die Eingabe von mathematischen Ausdrücken wesentlich vereinfachen. Die Paletten enthalten unter anderem griechische Symbole und mathematische Konstrukte wie Integrale, Begrenzungen, Summen, Quadratwurzeln und Potenzen. Neue, kontextabhängige Menüs

Menüs, die mit der rechten Maustaste aufgerufen werden, führen den Anwender durch den umfangreichen mathematischen Befehlsvorrat und erleichtern Formatierungen und Grafikbearbeitung. Der Anwender klickt einfach auf einen mathematischen Ausdruck oder eine Grafik und Maple V generiert hierfür ein spezifisches Menü an Befehlen.

Diese Leistungsmerkmale und eine Reihe von mathematische Highlights sind in Maple V Release 5.0 integriert, um die Benutzerfreundlichkeit und Flexibilität eines der weltweit führenden Computer Algebra Systeme nochmals zu verbessern.

Maple V Release 5.0 steht für alle gängigen Plattformen zur Verfügung. Dazu zählen die Betriebssysteme Windows 95, Windows NT, Windows 3.1+, Macintosh, Power Macintosh sowie alle gängigen UNIX-Systeme inklusive Linux. Anmerkung: die Funktionalität ist abhängig vom Betriebsystem.

Andreas Himmeldorf (Aachen)

Berichte über Arbeitsgruppen

Computeralgebra am Lehrstuhl D für Mathematik der RWTH Aachen

Mitglieder der Arbeitsgruppe: Prof. Dr. G. Hiß, Prof. em Dr. J. Neubüser, Prof. Dr. H. Pahlings, Prof. Dr. U. Schoenwaelder, Dr. V. Felsch, Dr. U. Klein, Dr. F. Lübeck, Dr. J. Müller, T. Breuer, T. Merkwitz, M. Neunhöffer, M. Szőke. Arbeitsgebiete: Darstellungstheorie endlicher Gruppen, rechnergestützte Gruppen- und Darstellungstheorie, Erweiterung vorhandener Computeralgebra-Systeme. Hier eine Liste laufender Projekte, in denen der Einsatz von Rechnern eine wichtige Rolle spielt:

• Entwicklung des Programmpakets CHEVIE [1] zum Rechnen in Coxetergruppen, zum Konstruieren und Manipulieren generischer Charaktertafeln von Iwahori-Hecke-Algebren und endlichen Gruppen vom Lie-Typ. Anwendung der Ergebnisse in der Darstellungstheorie dieser Algebren und Gruppen.

• Konstruktion modularer Charaktertafeln sporadischer Gruppen als Beitrag zu einem zweiten Band des Modularen Atlas [2]. Hierzu werden neben GAP [3] die Programmpakete Meat-Axe und MOC benutzt und weiterentwickelt, insbesondere die Kondensationsmethoden der Meat-Axe. Anwendung der Ergebnisse etwa im Projekt zur Matrixgruppen-Erkennung.

• Strukturanalyse von Modulkategorien verschiedener Algebren. Hierzu werden die Meat-Axe mit neueren Erweiterungen sowie GAP verwendet. Insbesondere Berechnung von Green-Korrespondenten und Beiträge zur Verifikation der Vermutungen von Alperin und Dade.

• Weiterentwicklung und Betreuung der darstellungstheoretischen Algorithmen und Datensammlungen von GAP, insbesondere laufendes Ergänzen der Bibliothek der Charakter- und Markentafeln endlicher Gruppen. Beantwortung von Benutzeranfragen zu GAP aus dem Bereich der Darstellungstheorie.

• Weiterentwicklung und Betreuung von Restklassenmethoden in GAP, Xgap, Einsatz von GAP in der Lehre.

Literatur: M. Geck, G. Hiss, F. Lübeck, G. Malle, and G. Pfeiffer, CHEVIE — A system for computing and processing generic character tables, AAECC 7 (1996), 175–210. C. Jansen, K. Lux, R. A. Parker, and R. A. Wilson, An Atlas of Brauer Characters, Oxford University Press, 1995. M. Schönert et al., GAP — Groups, Algorithms, and Programming, Lehrstuhl D für Mathematik, RWTH Aachen, Germany, fifth ed., 1995.

Computeralgebra in der Lehre: Regelmäßige Vorlesungen über Computeralgebra und Kurse in Maple. Zusammenarbeit: Im Rahmen der oben genannten Projekte besteht eine dauernde aktive Zusammenarbeit mit den folgenden Arbeitsgruppen: Prof. Dr. W. Plesken (Lehrstuhl B für Mathematik, RWTH Aachen), Prof. Dr. B. H. Matzat (IWR Heidelberg), Prof. Dr. M. Broué (Équipe des groupes finis, Paris), Prof. Dr. E. Robertson, Dr. S. Linton (The GAP-team, St. Andrews), Prof. Dr. T. Shoji (Tokio), Prof. Dr. K. Lux (Tucson, Arizona), Prof. Dr. Á. Seress (Columbus, Ohio), Dr. D. Holt (Warwick), Prof. Dr. C. Wright (Oregon).

Weitere Informationen unter der URL: http://www.Math.RWTH-Aachen.De/LDFM

G. Hiß (Aachen)

The FRISCO Project

FRISCO: A Framework for Integrating Symbolic/Numeric Computation is an Esprit Long Term Research project which began in March 1996. The project lasts three years and is dedicated to developing state-of-the-art algorithms for solving systems of polynomial equations, and a framework for delivering implementations of those algorithms in an industrial environment. The FRISCO consortium consists of the Numerical Algorithms Group Ltd (NAG) as co-ordinating partner, with INRIA Sophia-Antipolis, IRMAR at the University of Rennes, the University of Cantabria and the University of Pisa as partners. This consortium brings together some of the main members of the PoSSo consortium, who under a previous Esprit project developed amongst other things a C++ library for solving polynomial systems, with some of the people directly involved in developing Aldor, the new compiler for the AXIOM computer algebra system.

A principle outcome of the first year of the project was a review of real-life industrial problems which can be reduced to systems of polynomials (this review can be found at http://www.nag.co.uk/ projects/FRISCO/reports/industry.ps). In the second year contacts with industry have continued and we have begun to tackle a number of problems in areas including Computer Aided Design, Engineering and Telecommunications. A lot of effort has been devoted to developing implementations of algorithms which are suitable for the solution of these kinds of problems. In the third year we expect to devote more effort to applying these algorithms in practice.

Until now the project has maintained two libraries of algorithmic code: the successor to the original PoSSo Library, written in C++; and the new BasicMath Library, written in Aldor. A key result of the second year is a mechanism to allow interoperability between programs written in Aldor and C++. This allows Aldor programmers to use highly efficient libraries such as those produced during the PoSSo project. Currently work is underway to allow the same kind of interaction between Aldor and Fortran. Further work has also been done on component software technologies to allow FRISCO algorithmic technology to be integrated into other applications via “black boxes” which use a simple interface to hide the complexity of the underlying mathematical algorithms from a non-expert user. During the third year the project hopes to develop a number of these black boxes for carrying out high-level tasks.

More information about FRISCO can be found at http://www.nag.co.uk/projects/FRISCO, or from the coordinator, miked@nag.co.uk.

Mike Dewar (Oxford)

Publikationen über Computeralgebra

• Bündgen, R., Termersetzungssysteme, Vieweg Verlagsges., ISBN 3-528-05652-5, 1998, pp. 228, DM 49,80.

• Blum, L., Cucker, F., Shub, M., Smale, S.,, Complexity and Real Computation, Springer Verlag, ISBN 0-387-98281-7, 1997, DM 79,00.

• Fröberg, Ralf, An Introduction to Gröbner Bases, Chichester, ISBN 0-471-97442-0, 1997, DM 95,00. Das Buch wird im nächsten Rundbrief besprochen.

• Grey A., Mezzino M., Pinsky M.A., Introduction to Ordinary Differential Equations with Mathematica – An Integrated Multimedia Approach, Springer Verlag, ISBN 0-387-94481-8, 1997, pp. 890. Das Buch wird in diesem Rundbrief auf Seite 20 besprochen.

• Klimek, G., Klimek, M., Discovering Curves and Surfaces with Maple, Springer Verlag, ISBN 0-387-94890-2, 1997, pp. 640, DM 86,00.

• Vasconcelos, Wolmer V., Computational Methods in Commutative Algebra and Algebraic Geometry, Springer Verlag, ISBN 0-540-60520-7, 1998, pp. 394, DM 118,00. Das Buch wird in diesem Rundbrief auf Seite 22 besprochen.

Besprechungen zu Büchern der Computeralgebra

• Grey A., Mezzino M., Pinsky M.A., Introduction to Ordinary Differential Equations with Mathematica – An Integrated Multimedia Approach

Verlag Springer: Berlin - Heidelberg - New York, ISBN 0-387-94481-8, 1997, 890 Seiten, CD-ROM.

Das vorliegende Buch stellt einen abgeschlossenen Kurs über gewöhnliche Differentialgleichungen dar, erweitert und multimedial angereichert durch den Einsatz des CAS Mathematica. Dass es sich im Wesentlichen um ein traditionelles Lehrbuch zum genannten Thema handelt, zeigt eine Zusammenfassung der behandelten Themen:

– Differentialgleichungen 1. Ordnung: separierbare, exakte, Bernoulli Differentialgleichungen. Existenz- und Eindeutigkeitsbeweise für deren Lösungen, Anwendungen dazu.

– Lineare Differentialgleichungen 2. Ordnung: Theorie zur Existenz und Eindeutigkeit der Lösungen, Wronski-Determinante, Reduktion der Ordnung, konstante Koeffizienten, Anwendungen.

– Lineare Differentialgleichungen höherer Ordnung: mit konstanten Koeffizienten, Variation der Parameter, seminumerische Lösungen.

– Numerische Lösungsmethoden: Euler, Heun, Runge-Kutta

– Lapace Transformation

– Systeme von linearen Differentialgleichungen: Existenz- und Eindeutigkeitssätze, Variation der Parameter, Lösung mit Hilfe der Laplace Transformation, Anwendungen.

– Stabilität und Anwendung von nichtlinearen Systemen

– Lösung von Differentialgleichungen 2. Ordnung mit Hilfe von Potenzreihen

– Frobenius Lösungen von Differentialgleichungen 2. Ordnung

Mathematica kommt auf zwei Arten zum Einsatz: Erstens ist der Text mit einer Unzahl von Mathematica-generierten Grafiken versehen, die der Visualisierung schwieriger Sachverhalte dienen und durchwegs ihren Zweck erfüllen. Zweitens haben die Autoren ein Zusatzpaket zu Mathematica geschrieben – ODE.m –, das eine erweiterte Behandlung von Differentialgleichungen erlaubt, als dies normal unter Mathematica möglich ist. So kann beim Aufruf von ODE z.B. die Art der Differentialgleichung und damit die Lösungsmethode angegeben werden. ODE.m läuft, obwohl es für Mathematica 2.2 geschrieben wurde auch unter Mathematica 3.x, neue Versionen können im Internet von

http://math.cl.uh.edu/ode/ode.html geladen werden.

Dem Vortragenden steht also ein das Stoffgebiet umfassend abdeckendes Buch zur Verfügung. Aber auch Studenten kann man dieses Buch empfehlen. Wie bereits erwähnt tragen die zahlreichen Grafiken durchaus zum Verständnis bei. Nach kurzer Einlernzeit in ODE, steht den Studenten ein mächtiges Werkzeug zur Verfügung, um Differentialgleichungen ohne den klassischen Ballast der seitenweise händischen Rechnung behandeln und lösen zu können. Sehr gut gelungen sind auch die Beispiele zu den einzelnen Abschnitten. Es gibt ca. 300 durchgerechnete, sowie weitere 650 nicht gelöste Beispiele, sowohl rein theoretischer, wie auch anwendungsorientierter Art. Zum Erlernern von Lösungsstrategien, wurden diese nach ihrer Erarbeitung in ”Procedure Boxes“ zusammengefasst, sodass, hat man den Typ einer Gleichung erkannt, das Abarbeiten des Lösungsalgorithmus’ leicht wird. Eine nette Idee sind auch die immer wieder vorkommenden historischen Anmerkungen mit Porträts der wichtigsten Mathematiker, die auf den entsprechenden Gebieten gearbeitet haben. Die beiliegende CD-ROM enthält das bereits erwähnte Paket ODE.m mit einem Reference Manual, Lösungsnotebooks zu den Übungen des Buches sowie einige Zusatznotebooks zu Themen, die im Buch nicht behandelt werden (z.B. Phasenporträt eines Pendels]. Alle Notebooks liegen sowohl im Mathematica 2.2 als auch im Mathematica 3.x Format vor, sodass es keine Versionsprobleme geben dürfte. Weiters gibt es einige mit Mathematica erzeugte Movies, die z.B. die Entwicklung des Lorentz-Attraktors zeigen, sowie auch Movies, die im Quick-Time Format vorliegen. Zusammenfassend kann man sagen, dass den Autoren ein Buch gelungen ist, das man sowohl Vortragenden als auch Studenten empfehlen kann, die einerseits Literatur über Differentialgleichung benötigen und zweitens einen ”halbklassischen“ (Klassisch mit Erweiterung durch den Einsatz eines CAS) zu diesem Thema suchen. Einziger kleiner Wermutstropfen: Der im Titel versprochene ”Multimedia Approach“ findet praktisch nur unvollständig statt, zu Multimedia gehört mehr als einige Movies.

Werner Cyrmon (Bad Fischau, Österreich)

• Klimek, G., Klimek, M., Discovering Curves and Surfaces with Maple

Springer-Verlag, Berlin-Heidelberg-New York, 1997, ISBN 0-387-94890-2, 120 Abb., 640 S., Hardcover DM 86,00.

Die Visualisierung von Ergebnissen spielt im wissenschaftlichen Rechnen eine immer größere Rolle. Dementsprechend verfügen die meisten Computeralgebrasysteme über immer ausgefeiltere und umfangreichere graphische Komponenten. Die meisten Benutzer können zwar ohne größere Probleme einfache Graphiken wie der Plot einer Funktion oder einer Punktmenge erstellen; das volle Potential der Systeme werden sie aber oft nicht einmal erahnen. Zur Erstellung komplexerer Graphiken, insbesondere dreidimensionaler Darstellungen von Flächen, fehlen häufig auch Grundkenntnisse bei so wichtigen Aspekten wie Perspektive oder Beleuchtung.

In dem vorliegenden Buch werden die graphischen Möglichkeiten von Maple detailliert besprochen. Es erläutert zum einen die wesentlichen Befehle mit ihren vielen Optionen sowie die dahinter stehenden Datenstrukturen. Zum anderen versucht es, dem Leser ein Basiswissen zu vermitteln. So werden Themen wie geometrische Transformationen, Schattenwurf, Perspektive oder Licht und Farbe behandelt. Dabei beschränkt sich das Buch — wie schon aus dem Titel klar hervorgeht — auf die Darstellung von Kurven und Flächen (inklusive Animationen).

Das Buch will kein umfassendes Lehrbuch der Computergraphik oder der Visualisierung sein. Wie die Autoren im Vorwort schreiben, soll es als ein ”pragmatic do-it-yourself guide“ dienen. Dieses Ziel haben sie auch sehr gut erreicht. Ein Ingenieur oder Naturwissenschaftler, der die graphische Darstellung seiner Ergebnisse verbessern möchte, aber auch jeder Benutzer von Maple, der einfach ein wenig mit der Graphik spielen will, wird in dem Buch viele nützliche Hinweise oder Anregungen finden.

Werner M. Seiler (Mannheim)

• Krawietz, A., Maple V für das Ingenieurstudium

Springer-Verlag, Berlin-Heidelberg-New York, 1997, ISBN 0-540-60223-2, ix+358 Seiten, DM 78,00.

Der Autor verfolgt mit seinem Buch zwei Ziele:

1. Die Darstellung wichtiger Maple-Elemente, die zur Lösung von Problemen der Ingenieurmathematik benötigt werden.

2. Die Vermittlung eines vertieften Verständnisses mathematischer Methoden und ihrer technischen Anwendungen.

Das Buch gliedert sich in die Kapitel 1. Grundkonzepte der Computeralgebra, 2. Funktionen, Kurven und Gleichungen,

3. Lineare Algebra, 4. Analysis, 5. Statistik und Wahrscheinlichkeitsrechnung, 6. Approximation von Funktionen, 7. Gewöhnliche Differentialgleichungen und 8. Partielle Differentialgleichungen und endet mit einem Literatur- sowie einem Sachwortverzeichnis. Viele Abschnitte enthalten Übungsvorschläge, jedoch ohne Lösungen. Dem Buch ist eine Diskette beigelegt, auf der die im Text beschriebenen interaktiven Maple-Sitzungen und -Prozeduren als DOS-Text-Dateien enthalten sind.

In Kapitel 1 behandelt der Autor Grundtechniken im Umgang mit Maple. Besprochen werden Variable, Listen, Mengen und Zahlen. Darüber hinaus werden in diesem Kapitel die Grundlagen der Maple-Programmierung (Kontrollstrukturen, Prozeduren) vorgestellt. Das zweite Kapitel ist den elementaren Funktionen und Gleichungen gewidmet. Eingegangen wird auf Funktionen zweier Variablen, dem Zeichnen von Kurven in Ebene und Raum sowie der Bestimmung der Lösungen von Gleichungen und Gleichungssystemen. Das Kapitel über lineare Algebra befaßt sich mit den Lösungen linearer Gleichungssysteme mit zwei- und drei Unbekannten. Als ausführliches Anwendungsbeispiel wird gezeigt, wie die Statik eines zweidimensionalen Fachwerks mit den entsprechenden Maple-Befehlen berechnet werden kann. Ein zweites Anwendungsbeispiel demonstriert die Berechnung von Flächenmomenten. Im vierten Kapitel werden Grenzwerte von Folgen und Funktionen sowie Taylorreihen, Ableitungen und Integrale untersucht. Das Kapitel zur Statistik und Wahrscheinlichkeitsrechnung befaßt sich mit Stichproben, linearer und nichtlinearer Regression, Fehlerschätzungen sowie diskreten Verteilungen. Als Anwendungsbeispiel wird die Auswertung eines physikalischen Versuchs (Schwingung eines Pendels) beschrieben. Im sechsten Kapitel werden Methoden erläutert, wie eine gegebene Funktion in Maple näherungsweise beschrieben werden kann. Zu diesem Zweck werden verschiedene Approximationsverfahren (z. B. Taylor-, Pade-, Polynom-Approximation, trigonometrische Approximation, Fourier-Reihen) vorgestellt. Ausführliche Anwendungsbeispiele enthalten die beiden Kapitel über Differentialgleichungen. Erwähnt seien an dieser Stelle die Wärmeleitung in Hohlzylindern, Schwingungen, Rand- und Eigenwertprobleme sowie ebene Potential- und Bipotentialprobleme.

Eine erfolgreiche Arbeit mit diesem Buch setzt gewisse Grundfertigkeiten voraus. So ist es beispielsweise notwendig, daß der Leser des Buchs die Oberfläche von Maple kennt und mit ihr umgehen kann. Das Buch kann und soll nicht als Handbuch oder Nachschlagewerk für Maple dienen. Ebenso wird vom Leser erwartet, daß er die behandelten mathematischen Teilgebiete sicher beherrscht.

Das erste der beiden obigen Ziele wird vom Autor durch eine große Anzahl schöner Anwendungsbeispiele erreicht. Diese Vielzahl von Beispielen ist es auch, die dieses Buch von einer Reihe anderer Werke zum Computeralgebrasystem Maple, die in den letzten Jahren erschienen sind, unterscheidet. Viele der Beispiele stammen aus der Physik oder der Technischen Mechanik, andere Gebiete (zum Beispiel Anwendungen aus der Elektrotechnik) kommen zu kurz oder fehlen vollständig.

Besonders geglückt ist der Abschnitt 7.2.4 über die Behandlung von Differentialgleichungen am Beispiel eines ”Doppelpendels“. An zehn numerischen Anfangswerten werden die möglichen Verhaltensmuster dargestellt. Mit solchen Beispielen vermag es der Autor daher durchaus, ein vertieftes Verständnis mathematischer Methoden zu vermitteln. Andererseits gibt es auch Abschnitte (z. B. 6.1 Taylor- und Pade-Approximation), die sich mehr oder weniger auf das Vorstellen von Maple-Befehlen beschränken und daher das zweite Ziel verfehlen.

Insgesamt ist das Buch durch die große Anzahl von Anwendungsbeispielen gewiß gelungen und kann Studenten der Ingenieurwissenschaften als Lektüre empfohlen werden. Es ist mit einem Preis von DM 78,– jedoch relativ teuer.

W. Struckmann (Braunschweig)

• Vasconcelos, W.V., Computational Methods in Commutative Algebra and Algebraic Geometry

Band 2 der Reihe Algorithms and Computations in Mathematics, Springer-Verlag Berlin-Heidelberg-New York 1998, ISBN 3-540-60520-7, pp. 394, DM 118.

Dieses Buch eines ausgewiesenen Experten auf dem Gebiet der konstruktiven Methoden in der kommutativen Algebra (um diese geht es im wesentlichen) ist ein weiterer Mosaikstein in der monographischen Aufarbeitung der Grundlagen des symbolischen Rechnens, der sich eine Reihe von Autoren in den letzten Jahren verstärkt gewidmet haben. Ein wichtiges Anliegen des Autors besteht darin, auf beschränktem Raum eine Vielzahl von Ideen aufzugreifen und in (oft allerdings nur dem fortgeschrittenen Leser verständlichen) Grundzügen darzustellen, die sich in den letzten Jahren bei quantitativen Untersuchungen im Bereich der kommutativen Algebra als fruchtbar erwiesen haben und bisher einzig über eine Vielzahl von Zeitschriftenartikel (wenn überhaupt) zugänglich waren. Um dieses Ziel zu erreichen, muß die Darlegung des umfangreichen begrifflichen Apparats der kommutativen Algebra rudimentär bleiben, was mit Hinblick auf existierende klassische (Kaplansky, Matsumura) oder neuere Texte ([E] oder [BH]) allerdings kein Mangel ist. Wissen im Umfang von [AM] wird vorausgesetzt, jedoch sollte der Leser auch einige Erfahrung mit den darüber hinaus in einem Anhang A (60 S.) “A Primer on Commutative Algebra” entwickelten Konzepten besitzen, die von Noetherschen Ringen über ganze Erweiterungen und Syzygientheorie bis hin zu lokaler Kohomologie und Liaison reichen. Kenntnisse über Hilbertreihen als einem der grundlegenden Werkzeuge für homogene Ideale sind ebenfalls nützlich, obwohl in einem von J.Herzog verfaßten Anhang B (24 S.) wichtige Ergebnisse über Hilbertfunktionen (G-Filtrierungen und die Rolle des Rees- sowie des assoziierten graduierten Rings; das Bayersche Deformationsargument; Hilbert-Funktionen, Auflösungen und lokale Kohomologie; Lex-Segment-Ideale, Bettizahlen und die Sätze von Green und Gotzmann) zusammenhängend entwickelt werden. Aus ähnlichem Grund wird auf eine fundierte Einführung in die Theorie der Gröbnerbasen verzichtet, obwohl diese das primum mobile eigentlich aller diskutierten konstruktiven Verfahren sind; es wird einfach davon ausgegangen, daß diese in einer genügend effizienten und umfassenden (Modulfall, Syzygienberechnung) Implementierung zur Verfügung stehen. Neben einer Zusammenstellung grundlegender Begriffe zu diesem Thema in Kapitel 1 (20 S.) wird deshalb im Anhang C (25 S., Autoren: D.Eisenbud, D.R.Grayson, M.E.Stillman) das Arbeiten mit der Version 2 des bekannten Spezialsystem Macaulay als einem für die entsprechenden Rechnungen bestens geeigneten Werkzeug vorgestellt.

Das Buch ist also kein Lehrbuch im herkömmlichen Verständnis, obwohl für Graduiertenseminare bei entsprechenden Vorkenntnissen durchaus geeignet (und vom Autor in Teilen auf verschiedenen Sommerschulen auch bereits eingesetzt). Es beginnt im wesentlichen dort, wo in den oben genannten grundlegenderen Monographien insbesondere konstruktive Aspekte nicht weiter vertieft werden und ist damit ein Mosaikstein im wirklichen Sinne. Es ist zugleich keine Monographie im herkömmlichen Verständnis, die sich an einem engen Thema erschöpfend abarbeitet, sondern mehr eine Fundgrube von Ideen und Ansätzen, in denen die verschiedensten Techniken der kommutativen Algebra bis hin zu tiefliegenden homologischen Methoden zusammenspielen. Sie sind, wie vom Autor gewohnt, oft unterschiedlich detailliert ausgearbeitet, beleuchten aber an vielen Stellen Querverbindungen, die man in dieser Kostellation in anderen Arbeiten selten findet. Für ein detaillierteres Studium insbesondere von Fragestellungen, die in den späteren Kapiteln aufgegriffen werden, wird der Leser deshalb kaum um die Konsultation der entsprechenden Zeitschriftenaufsätze herumkommen, sofern er sie nicht sowieso schon kennt. Als guter Leitfaden für ein solches Unterfangen ist das vorliegende Buch allerdings bestens geeignet und deshalb jedem, der sich ernsthaft mit konstruktiven Methoden in der kommutativen Algebra beschäftigen möchte, zu empfehlen. Um dem Leser dieser Rezension auch einen gewissen inhaltlichen Überblick zu ermöglichen, seien zum Abschluß aus den einzelnen Kapiteln die wichtigsten Themen aufgelistet, die aufgegriffen und auf konstruktive Aspekte hin untersucht werden: Kap. 2 (34 S.): Toolkit Endomorphismenringe und reguläre Elemente; Noether-Normalisierung und Noether-Komplexität; Fitting-Ideale; ganze Erweiterungen; Flachheit, generische Flachheit und Torsionsfreiheit; Cohen-Macaulay-Algebren und deren Hilbertfunktion. Kap. 3 (36 S.): Primärzerlegung Grundlegende Begriffe; Primärzerlegung von Monomidealen; äquidimensionale Zerlegung und homologische Methoden; grobe äquidimensionale Zerlegung ohne Exts; Lokalisierungstechniken und verschiedene Reduktionen auf Dimension 0; Primalitätstests für Ideale ohne Faktorisierung; Symbolische Potenzen. Kap. 4 (24 S.): Rechnen in Artinschen Algebren Artinsche Algebren und lineare Algebra, Sockel und Jacobson-Radikal; Radikal nulldimensionaler Ideale; Berechnung von Dekompositionen als verallgemeinerte Faktorisierung. Kap. 5 (22 S.): Nullstellensätze Radikalberechnung und Jacobi-Ideal; generische Sockel-Formeln; Konstruktion regulärer Sequenzen; obere Jacobi-Ideale und das Top-Radikal. Kap. 6 (38 S.): Ganzer Abschluß Multiplikationsring Hom_{R}(I, I); S_{2}-ifizierung; Desingularisierung in Kodimension 1; ganzer Abschluß eines Ideals und eines Rings. Kap. 7 (28 S.): Idealtransformierte und Invariantenringe Gleichungen für verschiedene Aufblasungsringe; Tangentialkegel und analytic spread; faktorieller Abschluß und symbolische Aufblasung; Unterringe von Polynomringen; Semigruppen-Ringe; SAGBI-Basen; Invariantenringe von linearen Gruppen.

Kap. 8 (8 S.): Zur Kohomologieberechnung über dem P^{n} (von D.Eisenbud) Kap. 9 (52 S.): Komplexitätsgrad eines graduierten Moduls Grad, arithmetischer und geometrischer Grad und Verallgemeinerungen; Schranke von Brownawell/Kollar im Nullstellensatz (o. Bew.); Reduktionszahl; arithmetischer Grad und Reduktionszahl; arithmetischer Grad und Erzeugendenzahl; Verallgemeinerungen des arithmetischen Grades, die verschiedene Homologien genauer berücksichtigen; verschiedene Schranken.

[AM] M.F.Atiyah, I.G.MacDonald: Introduction to Commutative Algebra, Addison-Wesley 1969

[BH] W.Bruns, J.Herzog: Cohen-Macaulay Rings, Cambridge Univ. Press 1993

[E] D.Eisenbud: Commutative Algebra with a View towards Algebraic Geometry, Springer 1994

Hans-Gert Gräbe (Leipzig)

Lehrveranstaltungen über Computeralgebra im SS 1998

• Rheinisch–Westfälische Technische Hochschule Aachen Einführungspraktikum in das Formelmanipulationssystem MAPLE, G. Hiß, U. Klein, V. Dietrich, P2 Praktikum Programmieren in MAPLE, G. Hiß, U. Klein, P4 Seminar über Maple, V. Dietrich, S2

• Technische Universität Berlin Einführung in die Codierungstheorie , M. Pohst, V4 Seminar Algorithmische Zahlentheorie und Algebra, M. Pohst, S2

• Technische Hochschule Darmstadt Seminar Lie-Gruppentheorie (Grundlagen, Berechnung der Erzeugenden und Anwendung in Mechanik, Physik und Technik), K.G. Roesner, S2

• Universität Dortmund Symbolisches und numerisches Lösen von Gleichungssystemen, H.M.Möller, V4 + Ü2 Oberseminar über alg. Geometrie und ihre Algorithmen, E. Becker, H.M.Möller, S2

• Universität Erlangen-Nürnberg Topics in Computer Algebra, V. Strehl, V2+Ü1

• Martin-Luther-Universität Halle(Saale)

Praktikum Mathematik mit Mathcad, H. Benker, P2 Mathematica-Begleitkurs zur Analysis, Benker, Carl, Kirsch, Pluschke, P2

• Universität Heidelberg Computeralgebra, G. Malle, V4 + Ü2

• Universität Kaiserslautern Computer Algebra, G. Pfister, V4 + Ü2 Praktkum Computeralgebra, G. Pfister, P4 Oberseminar Singularitätentheorie und Computeralgebra, G.-M. Greuel, G. Pfister S2

• Pädagogische Hochschule Karlsruhe Informatik II (unter Verwendung von Computeralgebra-Systemen), J. Ziegenbalg, V2 Hauptseminar Algorithmen, J. Ziegenbalg, HS2

• Universität Leipzig Kodierungstheorie, J. Apel, V2 Grundlegende Algorithmen der Computeralgebra, H.-G. Gräbe, V2 Algebraische Komplexitätstheorie , H.-G. Gräbe, V2 Mechanisiertes Theorembeweisen in der Geometrie , H.-G. Gräbe, V2 • Universität Linz, Research Institute for Symbolic Computation

Computer-Algebra für Fortgeschrittene, F. Winkler, V2 Geometrische Grundlagen für Symbolic Computation, S. Stifter, V2 + Ü1 Algorithmische Algebraische Geometrie, F. Winkler, V2 Mathematikunterricht mit DERIVE, B. Kutzler, V2 Elimination Theory, D. Wang, V2 Programming in MATHEMATICA, W. Windsteiger, V2 Projektseminar Computer Algebra, F. Winkler, S2

• Technische Universität München Computeralgebra II, M. Kaplan, V4

• Universität Oldenburg Gröbnerbasen und Anwendungen, W.Schmale, V4,Ü2 Seminar Ganze Zahlen, Polynome und Matrizen, W.Schmale S2

• Universität Passau Algebra und Logik, V. Weispfenning, V3 + Ü2 Oberseminar Computeralgebra, V. Weispfenning, S2

• Universität Rostock Symbolisches Rechnen, K. Hantzschmann, V2 Seminar Symbolisches Rechnen, A. Widiger, S2

• Universität Tübingen Praktikum Einführung in die Computeralgebra, R. Loos, P1 + Ü2 Seminar Generic Programming (in Englisch), R. Loos, D. Musser, S2

• Universität Ulm Symmetrie Analyse von Differentialgleichungen mit MathLie, G. Baumann, V2 Praktikum Symbolisches und Algebraisches Rechnen, ¿ G. Baumann, P4 Seminar zu Mathematische Methoden der Physik mit Mathematica , G. Baumann and T.F. Nonnenacher S2

• ETH Zürich Computer Algebra II, G. Gonnet, V2 + U1

Kurze Mitteilungen

• Lehrerfortbildung Computeralgebra an der Universität Heidelberg

Die Universität Heidelberg hat in Zusammanrbeit mit dem Oberschulamt Karlsruhe und Gymnasien im Fachbereich Mathematik am 14. Januar 1998 in den Räumen des IWR (Interdisziplinäres Zentrum für Wissenschaftliches Rechnen) eine Lehrerfortbildung zum Thema ’Computeralgebra’ veranstaltet. Diese stieß auf reges Interesse, ungefähr 75 Gymnasiallehrer aus der näheren und weiteren Umgebung Heidelbergs nahmen daran teil. Das Programm des Einführungstages bestand in je einem Übersichtsvortrag über ’Themen und Schwerpunkte der Computeralgebra’ von J. Grabmeier (IBM Heidelberg) sowie über ’Computeralgebra und Kryptographie’ von B. H. Matzat (Univ. Heidelberg). Diese Vorträge wurden am 22.1. bzw. 28.1. durch praktische Übungen am Rechner (vor allem mit MAPLE) ergänzt. Die Computeralgebra-Praktika wurden von R. Busam (Univ. Heidelberg) mit Unterstützung von M. Kuß geleitet.

• Preis zur Förderung der Qualität der Lehre der G. A. Müller-Stiftung 1997

Prof. Dr. Robert Keßler und Prof. Dr. Thomas Westermann, beide aus dem Fachbereich Naturwissenschaften der Fachhochschule Karlsruhe - Hochschule für Technik, wurden am 4. Dezember 1997 zu gleichen Teilen mit dem Preis der G. A. Müller-Stiftung zur Förderung der Qualität der Lehre ausgezeichnet. Der Preis wurde verliehen zur Anerkennung der persönlichen Lehrqualifikation, der innovativen Konzepten des Einsatzes von Computern in der Physik- bzw. Mathematikausbildung sowie für die zahlreichen Projektarbeiten. Den Preis, verbunden mit einer Prämie von 5.000 Mark, überreichte Heinz Heiler, Präsident der Landesvereinigung Bauwirtschaft Baden-Württemberg und Geschäftsführer der G. A. Müller GmbH & Co. KG, im Anschluß an den Vortrag von Prof. Dr. Westermann ”Mathematik alive - über den Einsatz von Computeralgebra-Systemen in der Mathematikvorlesung”.